### v1

# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

Abstract
Mixed effects models are widely used for modeling data with hierarchically grouped structures and high-cardinality categorical predictor variables. However, for high-dimensional crossed random effects, current standard computations relying on Cholesky decompositions can become prohibitively slow. In this work, we present novel Krylov subspace-based methods that address several existing computational bottlenecks. Among other things, we theoretically analyze and empirically evaluate various preconditioners for the conjugate gradient and stochastic Lanczos quadrature methods, derive new convergence results, and develop computationally efficient methods for calculating predictive variances. Extensive experiments using simulated and real-world data sets show that our proposed methods scale much better than Cholesky-based computations, for instance, achieving a runtime reduction of approximately two orders of magnitudes for both estimation and prediction. Moreover, our software implementation is up to 10’000 times faster and more stable than state-of-the-art implementations such as lme4 and glmmTMB when using default settings. Our methods are implemented in the free C++ software library GPBoost with high-level Python and R packages.
1 Introduction
Mixed effects models are widely used in various scientific disciplines for modeling data with hierarchically grouped structures (Laird and Ware, 1982; Pinheiro and Bates, 2000). High-dimensional crossed random effects occur frequently in practice, e.g., when modeling ratings in recommender systems with categorical grouping variables corresponding to customers and products (Gao and Owen, 2017; Ghosh et al., 2022a; Simchoni and Rosset, 2023) or, in general, when there are high-cardinality categorical predictor variables (Simchoni and Rosset, 2021; Sigrist, 2023b). However, current standard Cholesky decomposition-based computations can become prohibitively slow for high-dimensional crossed random effects. For instance, the cost of computing Cholesky factors of sparse matrices can be of cubic complexity in the dimension of the random effects (Pandolfi et al., 2024).
Various methods have been proposed to overcome this computational bottleneck, including method of moments estimation (Gao and Owen, 2017, 2020), collapsed Gibbs samplers (Papaspiliopoulos et al., 2020, 2023), backfitting algorithms (Ghosh et al., 2022b, a), composite likelihoods (Bellio et al., 2023; Xu et al., 2023), and variational inference (Xu et al., 2023; Goplerud et al., 2025). Furthermore, the preconditioned conjugate gradient (CG) method (Saad, 2003) with relatively simple diagonal or block-diagonal preconditioners has been applied to solve high-dimensional mixed model equations (Strandén and Lidauer, 1999, 2001; Tsuruta et al., 2001; Taskinen et al., 2017; Garrick et al., 2019; Vandenplas et al., 2018). Recently, Pandolfi et al. (2024) have used the CG method in Gibbs sampling for Bayesian generalized linear mixed models (GLMMs). For Gaussian likelihoods and the special case where all random effects have the same variance, Border and Becker (2019) and Cheng et al. (2023) use stochastic Lanczos quadrature (SLQ) (Ubaru et al., 2017) without preconditioning in combination with a shift-invariance property to calculate log-determinants in log-likelihoods. However, without preconditioning, SLQ-approximated log-determinants have high variance and are thus inaccurate; see our theoretical and empirical analyses below. Overall, there are two major unsolved computational challenges: (i) efficiently computing log-determinants in log-likelihoods and their gradients, and (ii) solving high-dimensional sparse linear systems with many right-hand sides for prediction.
In this article, we show how Krylov subspace methods, such as the preconditioned CG and SLQ methods, can be used for computationally efficient parameter estimation and the calculation of posterior predictive distributions for generalized mixed effects models (GMMs), a term we use to denote mixed effects models with linear and non-linear fixed effects functions and general response variable distributions. For non-Gaussian likelihoods, marginal likelihoods are approximated using the Laplace approximation since it is computationally efficient and converges asymptotically to the correct quantity. Supporting this argument, we show in Figure 1 the estimated variance parameter of one random effect in a model with two crossed random effects obtained with a Laplace approximation for varying numbers of repeated observations $d$ per random effect realization on simulated data with a Bernoulli likelihood and the setting described in Section 5.1 with $m=4^{\prime}000$. Estimation is done using the Krylov subspace methods introduced in this paper and repeated on $100$ simulated data sets for every $d$. Figure 1 shows that the variance parameter estimator has a small downward bias which vanishes as $d$ grows.
Figure 1: Estimated variance parameter $\sigma^{2}_{1}$ obtained with a Laplace approximation vs. different numbers of repeated observations $d$ per random effect realization for binary data. The red rhombi represent means and the whiskers are ±2$\pm$ 2$\times$ 2 × standard errors. The dashed line indicates the true parameter $\sigma^{2}_{1}=0.25$.
Krylov subspace methods allow for fast computations as they rely on matrix-vector multiplications with sparse matrices which can be trivially parallelized. We use the preconditioned CG and SLQ methods to solve sparse linear systems and to calculate log-determinants in log-likelihoods, respectively. Gradients of log-marginal likelihoods are calculated using stochastic trace estimation (STE) with almost no computational overhead once likelihoods are evaluated. We theoretically and empirically analyze several preconditioners with regard to the accuracy of SLQ approximations and the CG method’s convergence speed. Our theoretical results show that approximations with the SLQ method are expected to be accurate and that the CG method converges rapidly. In addition, we provide theoretical and empirical results showing that the symmetric successive over-relaxation (SSOR) preconditioner is superior compared to the popular diagonal preconditioner for both the CG and the SLQ methods. Furthermore, we introduce efficient and accurate simulation-based methods for calculating predictive variances when the number of prediction points is large. In experiments with simulated and real-world data, we obtain a reduction in runtime of approximately two orders of magnitude compared to Cholesky-based computations for estimation and prediction while having essentially identical accuracy. Moreover, our software implementation in the GPBoost library is up to 10’000 times faster and more stable than the state-of-the-art lme4 and glmmTMB packages when using default settings.
2 Preliminaries on Generalized Mixed Effects Models
We assume that the response variable $y=(y_{1},\dots,y_{n})^{T}\in\mathbb{R}^{n}$ follows a distribution with a density $p(y|\mu,\xi)=\prod_{i=1}^{n}p(y_{i}|\mu_{i},\xi)$, where $\mu\in\mathbb{R}^{n}$ are, potentially link function-transformed, parameters, and $\xi\in\Xi\subset\mathbb{R}^{r}$ are auxiliary parameters. For instance, $\mu_{i}$ and $\xi$ can be the mean and variance of a Gaussian distribution, the log-mean and the shape parameter of a gamma likelihood, or $\mu_{i}$ can be the logit-transformed success probability of a Bernoulli distribution for which there is no additional parameter $\xi$. The main parameters of interest $\mu$ are modeled as the sum of fixed $F(X)$ and random effects $Zb$:
$\mu=F(X)+Zb,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ b\sim N(0,\Sigma),$
where $X\in\mathbb{R}^{n\times p}$ contains predictor variables. In generalized linear mixed effects models (GLMMs), the fixed effects function is linear, $F(X)=X\beta$, but $F(\cdot)$ can also be modeled using machine learning methods such as regression trees (Hajjem et al., 2011; Sela and Simonoff, 2012; Fu and Simonoff, 2015), random forests (Hajjem et al., 2014), tree-boosting (Sigrist, 2022, 2023a), or deep neural networks (Simchoni and Rosset, 2021, 2023; Avanzi et al., 2024). The latent Gaussian variables $b=(b_{1}^{T},b_{2}^{T},\dots,b_{K}^{T})^{T}\in\mathbb{R}^{m}$ consist of $K$ grouped random effects $b_{j}\in\mathbb{R}^{m_{j}}$, $j\in\{1,\dots,K\}$, $m=\sum_{j=1}^{K}m_{j}$, and the covariance matrix $\Sigma\in\mathbb{R}^{m\times m}$ is diagonal and depends on a set of variance parameters $\theta=\{\sigma_{1}^{2},\sigma_{2}^{2},\dots,\sigma_{K}^{2}\}$:
$\Sigma=\text{diag}(\underbrace{\sigma_{1}^{2},\dots,\sigma_{1}^{2}}_{m_{1}-\text{times}},\underbrace{\sigma_{2}^{2},\dots,\sigma_{2}^{2}}_{m_{2}-\text{times}},\dots,\underbrace{\sigma_{K}^{2},\dots,\sigma_{K}^{2}}_{m_{K}-\text{times}})\in\mathbb{R}^{m\times m}.$
In addition, $Z=(Z_{1},Z_{2},\dots,Z_{K})\in\mathbb{R}^{n\times m}$, $Z_{j}\in\mathbb{R}^{n\times m_{j}}$, is usually a binary incidence matrix that maps the random effects to the corresponding observations, but it can also contain predictor variables when modeling random coefficients. We denote such a model as a generalized mixed effects model (GMM). Since we use the Laplace approximation, we further assume that $p(y|\mu,\xi)$ is log-concave in $\mu$.
2.1 Parameter estimation
[…]
$L^{LA}(y|F,\theta,\xi)=-\log p(y|\mu^{*},\xi)+\frac{1}{2}{b^{*}}^{T}\Sigma^{-1}b^{*}+\frac{1}{2}\log\det\left(\Sigma\right)+\frac{1}{2}\log\det\left(\Sigma^{-1}+Z^{T}WZ\right),$ (1)
where $\mu^{*}=F(X)+Zb^{*}$, $b^{*}=\operatorname*{argmax}_{b}\log p(y|\mu,\xi)-\frac{1}{2}b^{T}\Sigma^{-1}b$ is the mode of $p(y|\mu,\xi)p(b|\theta)$, and $W\in\mathbb{R}^{n\times n}$ is diagonal with $W_{ii}=-\frac{\partial^{2}\log p(y_{i}|\mu_{i},\xi)}{\partial\mu_{i}^{2}}\Big{|}_{\mu=\mu^{*}}$. The mode $b^{*}$ is typically found with Newton’s method (Williams and Rasmussen, 2006), and one iteration is given by
$b^{*t+1}=b^{*t}+\left(Z^{T}WZ+\Sigma^{-1}\right)^{-1}\left(Z^{T}\frac{\partial\log p(y|\mu^{*t},\xi)}{\partial b}-\Sigma^{-1}b^{*t}\right),\leavevmode\nobreak\ \leavevmode\nobreak\ t=0,1,\dots$ (2)
If a first- or second-order optimization method is used for minimizing $L^{LA}(y|F,\theta,\xi)$, gradients with respect to $\theta$, $F$, and $\xi$ are needed. These gradients can be found, e.g., in Sigrist (2023a). Note that gradients with respect to $F$ are used, for instance, for GLMMs since $\frac{\partial L^{LA}(y|F,\theta,\xi)}{\partial\beta}=X^{T}\frac{\partial L^{LA}(y|F,\theta,\xi)}{\partial F}$ when $F(X)=X\beta$.
For Gaussian likelihoods, we can write
$y=F(X)+Zb+\epsilon,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \epsilon\sim N(0,W^{-1}),\leavevmode\nobreak\ \leavevmode\nobreak\ W=\frac{1}{\sigma^{2}}I_{n},$
where $\sigma^{2}$ is the error variance and $I_{n}\in\mathbb{R}^{n\times n}$ an identity matrix. In this case, the Laplace approximation is exact, and the negative log-marginal likelihood can be equivalently calculated without determining the mode $b^{*}$ as follows:
$L(y|F,\theta,\xi)=\frac{n}{2}\log(2\pi)+\frac{1}{2}\log\det(\Psi)+\frac{1}{2}(y-F(X))^{T}\Psi^{-1}(y-F(X)),\leavevmode\nobreak\ \leavevmode\nobreak\ \Psi=Z\Sigma Z^{T}+W^{-1}.$ (3)
Since the dimension of the random effects $m$ is usually smaller than the number of observations $n$, one uses the Woodbury identity
$\Psi^{-1}=W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W$ (4)
for solving linear systems and the matrix determinant lemma to calculate
$\log\det(\Psi)=\log\det(\Sigma^{-1}+Z^{T}WZ)+\log\det(\Sigma)+\log\det(W^{-1}).$ (5)
Gradients, e.g., with respect to the variance parameters are given by
$\frac{\partial L(y|F,\theta,\xi)}{\partial\theta_{k}}=\frac{1}{2}\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial\Sigma^{-1}}{\partial\theta_{k}}+\Sigma^{-1}\frac{\partial\Sigma}{\partial\theta_{k}}\right)-\frac{1}{2}(y-F(X))^{T}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}(y-F(X)).$ (6)
The Fisher information $\mathcal{I}\in\mathbb{R}^{K\times K}$ for $\theta$, which can be used for Fisher scoring and for obtaining asymptotic confidence sets, is given by
[…]
2.2 Prediction
Consider the goal of making predictions for the latent variables
$\mu_{p}=F(X_{p})+Z_{po}b+Z_{pp}b_{p}\in\mathbb{R}^{n_{p}},$ (8)
where $X_{p}\in\mathbb{R}^{n_{p}\times p}$ is the predictor variable matrix of the predictions, $b_{p}\in\mathbb{R}^{m_{p}}$ are new random effects which are not in the training data random effects $b\in\mathbb{R}^{m}$, the matrix $Z_{po}\in\mathbb{R}^{n_{p}\times m}$ relates $b$ to $\mu_{p}$, and the matrix $Z_{pp}\in\mathbb{R}^{n_{p}\times m_{p}}$ relates the new random effects $b_{p}$ to $\mu_{p}$. We have
$\begin{split}\begin{pmatrix}b\\ \mu_{p}\end{pmatrix}&=\begin{pmatrix}0\\ F(X_{p})\end{pmatrix}+\begin{pmatrix}(I_{m},0_{m\times m_{p}})\\ (Z_{po},Z_{pp})\end{pmatrix}\begin{pmatrix}b\\ b_{p}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}0\\ F(X_{p})\end{pmatrix},\begin{pmatrix}\Sigma&\Sigma Z_{po}^{T}\\ Z_{po}\Sigma&Z_{po}\Sigma Z_{po}^{T}+Z_{pp}\Sigma_{p}Z_{pp}^{T}\end{pmatrix}\right),\end{split}$ (9)
where $(I_{m},0_{m\times m_{p}})\in\mathbb{R}^{m\times(m+m_{p})}$, $I_{m}\in\mathbb{R}^{m\times m}$ is an identity matrix, $0_{m\times m_{p}}\in\mathbb{R}^{m\times m_{p}}$ is a matrix of zeros, and $\Sigma_{p}=\text{Cov}(b_{p})$. Predictions for $\mu_{p}$ are obtained using the posterior predictive distribution $p(\mu_{p}|y,\theta,\xi)$ which is given by
[…]
where
$\omega_{p}$ $=F(X_{p})+Z_{po}b^{*},$ (10)
$\Omega_{p}$ $=Z_{po}\Sigma Z_{po}^{T}+Z_{pp}\Sigma_{p}Z_{pp}^{T}-Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T}$ (11)
$=Z_{pp}\Sigma_{p}Z_{pp}^{T}+Z_{po}\left(\Sigma^{-1}+Z^{T}WZ\right)^{-1}Z_{po}^{T}.$ (12)
This can be derived using the fact that a Laplace approximation for $p(y|F,\theta,\xi)$ is equivalent to the approximation $p(b|y,\theta,\xi)\approx\mathcal{N}\left(b^{*},\left(\Sigma^{-1}+Z^{T}WZ\right)^{-1}\right)$ since $p(y|F,\theta,\xi)=p(y|F,b,\xi)p(b|\theta)/p(b|y,\theta,\xi)$, and by applying standard results for conditional distributions of multivariate Gaussian distributions.
For the posterior predictive distribution $p(y_{p}|y,\theta,\xi)$ of the observable response variables $y_{p}$, $y_{p}|\mu_{p}\sim p(y_{p}|\mu_{p},\xi)$, the integral
[…]
must be additionally calculated. This is analytically intractable for most likelihoods but can be approximated using numerical integration or by simulating from $p(\mu_{p}|y,\theta,\xi)\approx\mathcal{N}\left(\omega_{p},\Omega_{p}\right)$.
For a Gaussian likelihood, the above predictive distributions are exact and available in closed form. The posterior mean in (10) can be equivalently calculated as
$\omega_{p}=F(X_{p})+Z_{po}\Sigma Z^{T}\Psi^{-1}(y-F(X)),$
and $p(y_{p}|y,\theta,\xi)$ is obtained from $p(\mu_{p}|y,\theta,\xi)$ by adding a diagonal matrix with the error variance to $\Omega_{p}$.
Note that, while the above predictive distributions might be less frequently used in classical statistical applications of GLMMs, they are important, e.g., for mixed effects machine learning models (Sigrist, 2022; Simchoni and Rosset, 2023) for high-cardinality categorical variables when probabilistic predictions are required.
3 Krylov Subspace Methods for Generalized Mixed Effects Models
Parameter estimation and prediction for models with high-dimensional crossed random effects involve several time-consuming operations. Specifically, there are the following bottlenecks all involving operations with the high-dimensional sparse matrix $\Sigma^{-1}+Z^{T}WZ\in\mathbb{R}^{m\times m}$. First, calculating linear solves $(\Sigma^{-1}+Z^{T}WZ)u=v$, $v\in\mathbb{R}^{m}$, (i) for quadratic forms of log-marginal likelihoods in (3) after applying the Woodbury identity given in (4), (ii) in Newton’s method for finding the mode, see (2), (iii) for implicit derivatives of the log-marginal likelihood, see Proposition 2.1 in Sigrist (2023a), and (iv) for predictive variances given in (12). The latter is particularly challenging as the number of prediction points $n_{p}$ is typically large, and the linear system in (12) thus contains many right-hand sides $Z_{po}^{T}$. Note that predictive variances are required not only for predictive distributions, but also for predictive means of non-Gaussian response variables, see (13). Second, calculating log-determinants $\log\det(\Sigma^{-1}+Z^{T}WZ)$ in (approximate) log-marginal likelihoods $L(y|F,\theta,\xi)$ given in (1) and (5) for non-Gaussian and Gaussian likelihoods, respectively. And third, calculating trace terms such as $\operatorname*{tr}((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}})$ for the derivatives of log-determinants and $\operatorname*{tr}(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}})$ for entries of the Fisher information. Traditionally, these operations are performed using a Cholesky decomposition of $\Sigma^{-1}+Z^{T}WZ$. In the following, we show how these operations can be done using Krylov subspace methods.
For linear solves with the matrix $\Sigma^{-1}+Z^{T}WZ$, we use the preconditioned conjugate gradient (CG) method, which solves a linear system $(\Sigma^{-1}+Z^{T}WZ)u=v$ by iteratively doing matrix-vector multiplications with $\Sigma^{-1}+Z^{T}WZ$. This can be done fast since $\Sigma^{-1}+Z^{T}WZ$ is sparse, and convergence is often achieved with $l\ll m$ iterations. Thus, linear solves can be calculated in $O(lm)$ time complexity. For completeness, the preconditioned CG algorithm is included in Appendix A.6.
Several techniques have been proposed in the literature to calculate log-determinants of large, symmetric positive definite matrices. Dong et al. (2017) find that stochastic Lanczos quadrature (SLQ) (Ubaru et al., 2017) achieves the highest accuracy and fastest runtime compared to other methods. In this article, we use the SLQ method in combination with the preconditioned CG method to approximate $\log\det(\Sigma^{-1}+Z^{T}WZ)$. Specifically, we first note that
$\log\det(\Sigma^{-1}+Z^{T}WZ)=\log\det(P)+\log\det(P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}),$ (14)
where $P$ is a symmetric positive definite preconditioner matrix; see Section 3.1 for more information on preconditioners. The last term is approximated with the SLQ method as follows:
$\log\det(P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}})\approx\frac{n}{t}\sum_{i=1}^{t}e_{1}^{T}\log(\tilde{T}_{i})e_{1},$ (15)
where $z_{1},\dots,z_{t}\in\mathbb{R}^{m}$ are i.i.d. random vectors with $\mathbb{E}[z_{i}]=0$ and $\mathbb{E}[z_{i}z_{i}^{T}]=P$, $\tilde{Q}_{i}\tilde{T}_{i}\tilde{Q}_{i}^{T}\approx P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}$ is a partial Lanczos decomposition obtained after $l$ steps of the Lanczos algorithm with $P^{-\frac{1}{2}}z_{i}/\|P^{-\frac{1}{2}}z_{i}\|_{2}$ as initial vector, $\tilde{Q}_{i}\in\mathbb{R}^{m\times l}$ has orthonormal columns, $\tilde{T}_{i}\in\mathbb{R}^{l\times l}$ is tridiagonal, and $e_{1}=(1,0,\dots,0)^{T}$. A derivation of this approximation can be found, e.g., in Kündig and Sigrist (2024, Appendix A.2). In this article, we use Gaussian random vectors $z_{i}\sim\mathcal{N}(0,P)$. Wenger et al. (2022) highlight that a decomposition with a preconditioner as in (14) leads to a reduction in variance. Intuitively, the more accurate the preconditioner $P\approx(\Sigma^{-1}+Z^{T}WZ)$, the smaller $\log\det(P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}})$, and thus the smaller the variance of its stochastic approximation.
As in Gardner et al. (2018), we use a technique from Saad (2003) to calculate the partial Lanczos tridiagonal matrices $\tilde{T}_{1},\dots,\tilde{T}_{t}$ from the coefficients of the preconditioned CG algorithm when solving $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{1},\dots,(W+\tilde{\Sigma}^{-1})^{-1}z_{t}$ $t$ times; see Appendix A.6. In doing so, we avoid running the Lanczos algorithm, which brings multiple advantages: Numerical instabilities are not an issue, storing $\tilde{Q}_{i}$ is not necessary, and the linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}$ can be reused in the stochastic trace estimation (STE) for calculating derivatives of the log-determinant, e.g., as follows:
$\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}\approx\frac{1}{t}\sum_{i=1}^{t}((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i})^{T}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}P^{-1}z_{i}.$
Gradients can thus be calculated with minimal computational overhead once the likelihood is calculated. In Appendix A.4, we show in detail how to calculate derivatives of log-determinants with respect to $\theta$, $F$, $b^{*}$, and $\xi$, using STE and a form of variance reduction with control variates based on the SSOR preconditioner introduced in the next section. Trace terms of the Fisher information in (7) can also be approximated with STE as follows:
$\begin{split}\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)=&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\left(\left(Z_{k}Z_{k}^{T}W-Z_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)z_{i}\right)^{T}\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(WZ_{l}Z_{l}^{T}-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}\right)z_{i},\end{split}$ (16)
where we use the Woodbury identity, $\frac{\partial\Psi}{\partial\theta_{k}}=Z_{k}Z_{k}^{T}$, and Gaussian random vectors $z_{i}\sim\mathcal{N}(0,I_{n})$. Linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}Wz_{i}$ and $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}z_{i}$ in (16) can be computed with the CG method. In experiments, we have found that an alternative STE for the Fisher information given in Appendix A.5 leads to less accurate estimates than the STE in (16) (results not tabulated).
3.1 Preconditioners
Preconditioners reduce the variance of approximations of log-determinants and their derivatives and accelerate the convergence speed of the CG method. To practically use a matrix $P$ as a preconditioner, we need to construct it, perform linear solves with it, calculate $\log\det(P)$, and sample from $\mathcal{N}(0,P)$ in a computationally efficient manner. In the following, we describe several preconditioners that we consider in this article.
3.1.1 SSOR preconditioner
The “symmetric successive over-relaxation” (SSOR) preconditioner is given by
$P_{\text{SSOR}}=(L+D)D^{-1}(L+D)^{T},$ (17)
where $D$ is a diagonal matrix with diagonal entries $(D)_{ii}=(\Sigma^{-1}+Z^{T}WZ)_{ii}$ and $L$ is a strictly lower-triangular matrix with entries $(L)_{ij}=\mathbf{1}_{\{i>j\}}(\Sigma^{-1}+Z^{T}WZ)_{ij}=\mathbf{1}_{\{i>j\}}(Z^{T}WZ)_{ij}$ such that $\Sigma^{-1}+Z^{T}WZ=L+L^{T}+D$.
3.1.2 ZIC preconditioner
The zero fill-in incomplete Cholesky (ZIC) factorization reported in Appendix A.7 gives the ZIC preconditioner $P_{\text{ZIC}}=LL^{T}\approx\Sigma^{-1}+Z^{T}WZ$, where $L\in\mathbb{R}^{m\times m}$ is a sparse lower triangular matrix that has the same sparsity pattern as $\Sigma^{-1}+Z^{T}WZ$. Unfortunately, we sometimes observe breakdowns (Scott and Tůma, 2014), i.e., (clearly) negative numbers in the calculations of square roots.
3.1.3 Diagonal preconditioner
The diagonal preconditioner $P_{\text{Diag}}$ is given by $P_{\text{Diag}}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})$. As mentioned in the introduction, this is currently the most popular preconditioner for the CG method for random effects models.
3.1.4 Other preconditioners
We additionally consider two low-rank preconditioners obtained by applying low-rank approximations to the matrix $Z^{T}WZ\approx L_{k}L_{k}^{T}$, $L_{k}\in\mathbb{R}^{m\times k}$ to obtain $P=\Sigma^{-1}+L_{k}L_{k}^{T}$. For these low-rank preconditioners, linear solves and log-determinants can be computed using the Woodbury matrix identity and the matrix determinant lemma, respectively, and the costs thus scale linearly with $m$. First, we apply the pivoted Cholesky decomposition (Harbrecht et al., 2012) to $Z^{T}WZ$ to obtain $L_{k}$ and denote the resulting preconditioner as $P_{\text{PChol}}$. This is a state-of-the-art approach for Gaussian Process regression (Gardner et al., 2018). Second, we apply a partial Lanczos algorithm to approximate $Z^{T}WZ\approx L_{k}L_{k}^{T}$ and denote the corresponding preconditioner as $P_{\text{Lanczos}}$.
3.2 Predictive variances
The computational bottleneck for prediction is the calculation of the posterior predictive variances $\text{diag}(\Omega_{p})$ in (12). These are required for predictive distributions of the latent and response variables as well as for predictive means of non-Gaussian response variables. Calculating $\text{diag}(\Omega_{p})$ requires solving $n_{p}$ linear equation systems with the matrix $\Sigma^{-1}+Z^{T}WZ$. For high-dimensional crossed random effects and large $n_{p}$, solving these linear systems with the CG method can thus be computationally prohibitive. We develop and compare three solutions for this: two simulation-based methods relying on the CG method and an approach based on the preconditioned Lanczos algorithm.
3.2.1 Predictive variances using a stochastic estimator for the diagonal of a matrix
Bekas et al. (2007) propose to approximate the diagonal of a matrix $A\in\mathbb{R}^{n_{p}\times n_{p}}$ as $\text{diag}(A)\approx\frac{1}{s}\sum_{i=1}^{s}z_{i}\odot Az_{i}$, where $\odot$ denotes the Hadamard product and $z_{i}\in\mathbb{R}^{n_{p}}$ are Rademacher random vectors with entries ±$\pm 1$. Algorithm 1 presents our method for calculating $\text{diag}(\Omega_{p})$ by stochastically approximating $\text{diag}(Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})$ in (12) based on the approach of Bekas et al. (2007). Note that Gyger et al. (2024) have used a similar algorithm for predictive covariances in Gaussian process regression. Linear solves for the stochastic estimator are computed with the preconditioned CG method. We additionally apply variance reduction in Algorithm 1 by using a control variate based on $\text{diag}(Z_{po}P^{-1}Z_{po}^{T})$:
$\begin{split}\text{diag}(Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})\approx&\;c\odot\text{diag}(Z_{po}P^{-1}Z_{po}^{T})\\ &+\frac{1}{t}\sum_{i=1}^{t}\underbrace{z_{i}\odot Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}z_{i}}_{\text{=:$h(z_{i})$}}-c\odot\underbrace{z_{i}\odot Z_{po}P^{-1}Z_{po}^{T}z_{i}}_{\text{=:$r(z_{i})$}},\end{split}$
where $c\in\mathbb{R}^{n_{p}}$ are the optimal weights for the variance reduction given by $c_{j}=\widehat{\text{Cov}}([h(z_{i})]_{j},[r(z_{i})]_{j})/\widehat{\text{Var}}([r(z_{i})]_{j})$, for $j=1,2,\dots,n_{p}$. This variance reduction is important as otherwise the variance of the stochastic approximation is much larger (results not tabulated). Algorithm 1 can be trivially parallelized and results in an unbiased and consistent approximation for $\text{diag}(\Omega_{p})$; see Appendix A.3 for a proof of Proposition 3.1. Assuming that $P^{-\frac{1}{2}}Z_{po}^{T}$ is precomputed, $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$, the computational complexity of the algorithm is $O(sn_{p}+slm+s\psi)$, where $l$ denotes the number of CG iterations, $s$ the number of simulation iterations, and $\psi$ the number of non-zero entries in $P^{-\frac{1}{2}}Z_{po}^{T}$.
Algorithm 1 Approximate predictive variances using stochastic estimator for the diagonal of a matrix
1:Matrices $Z_{po},Z,\Sigma,Z_{pp},\Sigma_{p}$, and $W$
2:Approximate predictive variances $\text{diag}(\hat{\Omega}_{p})$
3:for $i\leftarrow 1$ to $s$ do
[…]
5:     $z_{i}^{(2)}\leftarrow Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}z_{i}^{(1)}$
6:     $z_{i}^{(3)}\leftarrow Z_{po}P^{-1}Z_{po}^{T}z_{i}^{(1)}$
7:end for
8:$c_{j}\leftarrow\widehat{\text{Cov}}((z_{i}^{(1)}\odot z_{i}^{(2)})_{j},(z_{i}^{(1)}\odot z_{i}^{(3)})_{j})/\widehat{\text{Var}}((z_{i}^{(1)}\odot z_{i}^{(3)})_{j})$, $j=1,2,\dots,n_{p}$, $c=(c_{1},\dots,c_{n_{p}})^{T}$
9:$\text{diag}(\hat{\Omega}_{p})\leftarrow\text{diag}(Z_{pp}\Sigma_{p}Z_{pp}^{T})+c\odot\text{diag}(Z_{po}P^{-1}Z_{po}^{T})+\frac{1}{s}\sum_{i=1}^{s}(z_{i}^{(1)}\odot z_{i}^{(2)}-c\odot z_{i}^{(1)}\odot z_{i}^{(3)})$
Proposition 3.1.
Algorithm 1 produces an unbiased and consistent estimator $\text{diag}(\hat{\Omega}_{p})$ for the predictive variances $\text{diag}(\Omega_{p})$ given in (12).
We have also considered a similar algorithm using the alternative version in (11) for $\Omega_{p}$ instead of (12). Specifically, $\text{diag}(Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T})$ is approximated stochastically with the approach of Bekas et al. (2007) using the Woodbury identity, and all other terms are computed deterministically. However, experiments show that such an algorithm results in less accurate estimates compared to Algorithm 1 (results not tabulated).
3.2.2 Predictive covariances using simulation
Adopting an approach proposed in Kündig and Sigrist (2024), predictive covariances $\Omega_{p}$ can be approximated using expression (12) and sampling from a Gaussian distribution with covariance matrix $Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}$. Algorithm 2 presents this approach. The linear solves $(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}^{(3)}$ can be done using the preconditioned CG method and the computational complexity of Algorithm 2 is thus $O(sn+slm+sn_{p})$. This algorithm can also be trivially parallelized and results in an unbiased and consistent approximation for $\Omega_{p}$; see Appendix A.3 for a proof of Proposition 3.2. In contrast to Algorithm 1, this method can be used to calculate the entire covariance matrix and not just the variances on the diagonal. Furthermore, Algorithm 2 can be adapted to compute only the predictive variances $\text{diag}(\boldsymbol{\Omega}_{p})$ by summing $\boldsymbol{z}_{i}^{(4)}\odot\boldsymbol{z}_{i}^{(4)}$ in Line 7. We have also considered applying variance reduction to the estimator in Algorithm 2 by using a control variate based on $Z_{po}P^{-1}Z_{po}^{T}$. However, we found that the variance reduction was very small for predictive variances since $\widehat{\text{Cov}}((z_{i}^{(4)}(z_{i}^{(4)})^{T})_{jk},(z_{i}^{(5)}(z_{i}^{(5)})^{T})_{jk})$, $z_{i}^{(5)}=Z_{po}P^{-\frac{T}{2}}z_{i}^{(1)}$ and $j,k=1,2,\dots,n_{p}$, is often close to zero.
Algorithm 2 Approximate predictive covariances using simulation
1:Matrices $Z_{po},Z,\Sigma,Z_{pp},\Sigma_{p}$, and $W$
2:Approximate predictive covariances $\hat{\Omega}_{p}$
3:for $i\leftarrow 1$ to $s$ do
[…]
5:     $z_{i}^{(4)}\leftarrow Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}^{(3)}$
6:end for
7:$\hat{\Omega}_{p}\leftarrow Z_{pp}\Sigma_{p}Z_{pp}^{T}+\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$
Proposition 3.2.
Algorithm 2 produces an unbiased and consistent estimator $\hat{\Omega}_{p}$ for the predictive covariances $\Omega_{p}$ given in (12).
Alternatively, we can use the expression in (11) and approximate $Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T}$ using simulation. Algorithm 3 presents this approach. The linear solve $\Psi^{-1}z_{i}^{(3)}$ can be done using the Woodbury identity and the preconditioned CG method. The computational complexity of Algorithm 3 is thus $O(sn+slm+sn_{p})$. This algorithm also results in an unbiased and consistent approximation for $\Omega_{p}$; see Appendix A.3 for a proof of Proposition 3.3.
Proposition 3.3.
Algorithm 3 produces an unbiased and consistent estimator $\hat{\Omega}_{p}$ for the predictive covariances $\Omega_{p}$ given in (11).
Algorithm 3 Approximate predictive covariances using simulation
1:Matrices $Z_{po},Z,\Sigma,Z_{pp},\Sigma_{p}$, and $W$
2:Approximate predictive covariances $\hat{\Omega}_{p}$
3:for $i\leftarrow 1$ to $s$ do
[…]
5:     $z_{i}^{(4)}\leftarrow Z_{po}\Sigma Z^{T}\Psi^{-1}z_{i}^{(3)}$
6:end for
7:$\hat{\Omega}_{p}\leftarrow Z_{po}\Sigma Z_{po}^{T}+Z_{pp}\Sigma_{p}Z_{pp}^{T}-\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$
Moreover, we can apply the Woodbury identity in (11) and obtain
$\begin{split}Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T}=&Z_{po}\Sigma Z^{T}WZ\Sigma Z_{po}^{T}-Z_{po}\Sigma Z^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}.\end{split}$ (18)
We have additionally considered an algorithm, where only the last term in (18) is approximated by simulation and all other terms in (11) are calculated deterministically. However, we observe in experiments that such an algorithm leads to less accurate estimates than when using Algorithm 2 and Algorithm 3 (results not tabulated).
3.2.3 Predictive variances using preconditioned Lanczos algorithms
Pleiss et al. (2018) use the Lanczos algorithm to approximate predictive covariance matrices for Gaussian process regression. We adapt and extend this approach for GMMs. Specifically, we approximate $\text{diag}(Z_{po}\Sigma Z^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T})$ in (18) by running the Lanczos algorithm with the matrix $\Sigma^{-1}+Z^{T}WZ$ using the normalized average of the column vectors of $Z^{T}WZ\Sigma Z_{po}^{T}$ as initial value and approximate $(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}\approx\tilde{Q}\tilde{T}^{-1}\tilde{Q}^{T}Z^{T}WZ\Sigma Z_{po}^{T}$, where $\tilde{Q}\tilde{T}\tilde{Q}^{T}\approx\Sigma^{-1}+Z^{T}WZ$ is obtained after $k$ steps of the Lanczos algorithm, $\tilde{Q}\in\mathbb{R}^{n\times k}$ has orthonormal columns, and $\tilde{T}\in\mathbb{R}^{k\times k}$ is tridiagonal. $\text{diag}(Z_{po}\Sigma Z^{T}WZ\Sigma Z_{po}^{T})$ and the other terms in (11) are deterministically computed, and predictive variances are thus approximated by
$\begin{split}\text{diag}(\hat{\Omega}_{p})\approx\;&\text{diag}(Z_{po}\Sigma Z_{po}^{T})+\text{diag}(Z_{pp}\Sigma_{p}Z_{pp}^{T})-\text{diag}(Z_{po}\Sigma Z^{T}WZ\Sigma Z_{po}^{T})\\ &+\text{diag}(Z_{po}\Sigma Z^{T}WZ\tilde{Q}\tilde{T}^{-1}\tilde{Q}^{T}Z^{T}WZ\Sigma Z_{po}^{T}).\end{split}$ (19)
The computation of a rank $k$ Lanczos approximation requires one matrix-vector multiplication with $\Sigma^{-1}+Z^{T}WZ$ in each iteration. Since the Lanczos algorithm suffers from numerical stability issues due to loss of orthogonality, we use a full reorthogonalization scheme (Wu and Simon, 2000). Computing the Lanczos approximation then has complexity $O(k^{2}m)$. Furthermore, we consider a similar way of preconditioning as in Kündig and Sigrist (2024) by applying the partial Lanczos algorithm to $P^{-\frac{1}{2}}(\Sigma^{-1}+Z^{T}WZ)P^{-\frac{T}{2}}\approx\tilde{Q}\tilde{T}\tilde{Q}$ and correcting for this by using $P^{-\frac{T}{2}}\tilde{Q}\tilde{T}^{-1}\tilde{Q}P^{-\frac{1}{2}}Z^{T}WZ\Sigma Z_{po}^{T}\approx(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ\Sigma Z_{po}^{T}$ as approximation. Note that the Lanczos algorithm requires a symmetric matrix and in contrast to the preconditioned CG method, we need to explicitly calculate a factor $P^{-\frac{1}{2}}$, where $P=P^{\frac{1}{2}}P^{\frac{T}{2}}$. This is possible for the diagonal, SSOR, and ZIC preconditioners. However, in our experiments, this form of preconditioning does not lead to more accurate predictive variance estimates (results not tabulated).
3.3 Software implementation
The methods presented in this article are implemented in the GPBoost library written in C++ with Python and R interfaces, see https://github.com/fabsig/GPBoost.***Krylov subspace methods can be enabled via the parameter matrix_inversion_method = "iterative" and the preconditioner is chosen via the parameter cg_preconditioner_type. For linear algebra calculations, we use the Eigen library version 3.4.99 and its sparse matrix algebra operations whenever possible. Multi-processor parallelization is done using OpenMP.
4 Convergence analysis
In the following, we theoretically analyze the properties of the Krylov subspace methods presented in this article. We denote by $\lambda_{\min}=\lambda_{m}\leq\dots\leq\lambda_{1}=\lambda_{\max}$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. The eigenvalues of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ determine the convergence speed and accuracy of the CG and SLQ methods. However, there are important difference among the two methods. It is known that the CG method converges fast if the majority of the eigenvalues are clustered together and an effective condition number $\kappa_{m-l,k}=\frac{\lambda_{k}}{\lambda_{m-l}}$ (Van der Sluis and van der Vorst, 1986) is small for small $k,l\geq 1$, despite the presence of a few outlying small and large eigenvalues; see, e.g., Van der Vorst (2003) and Nishimura and Suchard (2022). In particular, a large condition number $\kappa=\frac{\lambda_{\max}}{\lambda_{\min}}$ is not necessarily a problem for the convergence speed of the CG method. Pandolfi et al. (2024) have recently shown that, for the diagonal preconditioner and a Gaussian likelihood, the majority of the eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ are clustered together except for a few small eigenvalues. This likely explains the relatively popular application of the diagonal preconditioner for the CG method in the prior literature (Tsuruta et al., 2001; Taskinen et al., 2017; Garrick et al., 2019; Pandolfi et al., 2024). In line with this, our experiments in Section 5 show that the CG method, whether used with the diagonal, SSOR, or ZIC preconditioner, or even without preconditioning, yields similar runtimes.
The situation is different for the SLQ method. Here, the smallest and largest eigenvalues, $\lambda_{\min}$ and $\lambda_{\max}$, are important for its convergence properties and accuracy as outlined in the following. In brief, the Lanczos algorithm factorizes a symmetric matrix as $A=QTQ^{T}$, where $Q\in\mathbb{R}^{n\times n}$ is orthonormal, and $T\in\mathbb{R}^{n\times n}$ is a tridiagonal matrix. This decomposition is computed iteratively, and $k$ iterations result in an approximation $A\approx\tilde{Q}\tilde{T}\tilde{Q}^{T}$, where $\tilde{Q}\in\mathbb{R}^{n\times k}$ contains the first $k$ columns of $Q$, and $\tilde{T}\in\mathbb{R}^{k\times k}$ the corresponding coefficients of $T$. Using $\log\det(A)=\operatorname*{tr}\log(A)$ and $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$, SLQ then applies stochastic trace estimation to approximate $\log\det(A)$; see (15), Ubaru et al. (2017), and Kündig and Sigrist (2024, Appendix A.2) for more details. The logarithm function used in $\log(A)\approx\tilde{Q}\log(\tilde{T})\tilde{Q}^{T}$ is mostly determined by the extremal eigenvalues of $A=P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. Furthermore, it is known that the Lanczos algorithm approximates the smallest eigenvalue $\lambda_{m}$ faster when the difference $\lambda_{1}-\lambda_{m}$ is smaller and when $\frac{\lambda_{m-1}-\lambda_{m}}{\lambda_{1}-\lambda_{m-1}}$ is larger (Golub and Van Loan, 2013). Similarly, the largest eigenvalue $\lambda_{1}$ is approximated the faster, the smaller $\lambda_{1}-\lambda_{m}$ and the larger $\frac{\lambda_{1}-\lambda_{2}}{\lambda_{2}-\lambda_{m}}$. Similar results can also be obtained for other eigenvalues. This means that the difference between the largest and smallest eigenvalues matters much more for the SLQ method compared to the CG method. In line with this, we find in our experiments in Section 5 that preconditioners for which the difference between the largest and smallest eigenvalues is small, such as the SSOR preconditioner, result in much more accurate log-determinant approximations compared to the diagonal preconditioner for which this difference is larger.
In addition to the above-mentioned intuitive and empirical arguments, there is convergence theory showing that the error of an SLQ-approximated log-determinant depends on the condition number $\kappa=\lambda_{1}/\lambda_{m}$ of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$. Below, we restate Theorem 3.2 of Kündig and Sigrist (2024) for our setting, which provides a stochastic error bound. We are not aware of any convergence results that relate an effective condition number $\kappa_{m-l,k}=\lambda_{k}/\lambda_{m-l}$ to the approximation error of an SLQ-approximated log-determinant.
Theorem 4.1.
Let $\kappa=\frac{\lambda_{1}}{\lambda_{m}}$ denote the condition number of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$, $C_{mt}=\frac{Q_{\chi^{2}_{mt}}({1-\eta/2})}{mt}$, where $Q_{\chi^{2}_{mt}}(\cdot)$ is the quantile function of a $\chi^{2}-$distribution with $mt$ degrees of freedom, and $\widehat{\Gamma}=\frac{1}{t}\sum_{i=1}^{t}\|P^{-\frac{1}{2}}z_{i}\|_{2}^{2}e_{1}^{T}\log(\tilde{T}_{i})e_{1}$ an SLQ approximation. If the SLQ method is run with $l\geq\frac{\sqrt{3\kappa}}{4}\log\left(\frac{C_{mt}20\log(2(\kappa+1))\sqrt{2\kappa+1}}{\epsilon}\right)$ preconditioned CG steps and $t\geq\frac{32}{\epsilon^{2}}\log(\kappa+1)^{2}\log\left(\frac{4}{\eta}\right)$ number of random vectors, the following holds for any preconditioner $P$:
$P(|\widehat{\Gamma}-\log\det(P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2})|\leq\epsilon m)\geq 1-\eta,$ (20)
where $\epsilon,\eta\in(0,1)$.
In the following, we analyze the extremal eigenvalues of the preconditioned matrices $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for different preconditioners $P$. We denote by $\lambda_{\min}^{\text{P}}=\lambda_{m}^{\text{P}}\leq\dots\leq\lambda_{1}^{\text{P}}=\lambda_{\max}^{\text{P}}$ the eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for a preconditioner $P$. In addition, the eigenvalues of the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ are denoted by $\lambda_{\min}^{\text{none}}=\lambda_{m}^{\text{none}}\leq\dots\leq\lambda_{1}^{\text{none}}=\lambda_{\max}^{\text{none}}$
4.1 SSOR preconditioner
Theorem 4.2 and Corollary 4.1 derive the extremal eigenvalues, as well as lower and upper bounds, of the SSOR-preconditioned matrix $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$, for general non-Gaussian and Gaussian likelihoods, respectively. In particular, Theorem 4.2 and Corollary 4.1 imply that the largest and smallest eigenvalues, $\lambda_{\max}^{\text{SSOR}}$ and $\lambda_{\min}^{\text{SSOR}}$, of $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$ do not depend directly on neither the sample size $n$ nor the dimension of the random effects $m$ but rather only on the number of repeated observations per random effect, denoted by $d=\frac{n}{m_{1}}=\frac{n}{m_{2}}$ when having a balanced design. Furthermore, the results show that the ratio $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ and the difference between the largest and smallest eigenvalues of $P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}$ are increasing in the number of repeated observations per random effect; see, e.g., (26).
Theorem 4.2.
Denote $D_{1}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{m_{1}\times m_{1}}$ and $D_{2}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{(m-m_{1})\times(m-m_{1})}$. If $K=2$, the following hold true for the SSOR preconditioner defined in Section 3.1.1:
$P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}=I_{m}-\begin{pmatrix}0&0\\ 0&D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}\end{pmatrix},$ (21)
$\lambda_{m_{1}}^{\text{SSOR}}=\dots=\lambda_{1}^{\text{SSOR}}=\lambda_{\max}^{\text{SSOR}}=1,$ (22)
$1-\frac{1}{(\min(\Sigma^{-1}_{ii}(Z^{T}WZ)_{ii}^{-1})+1)^{2}}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\frac{1}{(\max(\Sigma^{-1}_{ii}(Z^{T}WZ)_{ii}^{-1})+1)^{2}},$ (23)
$1-\max((D_{1}^{-1})_{ii})\max((D_{2}^{-1})_{ii})\max((Z^{T}WZ)_{ii})^{2}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\min((D_{1}^{-1})_{ii})\min((D_{2}^{-1})_{ii})\min((Z^{T}WZ)_{ii})^{2}.$ (24)
A proof of Theorem 4.2 can be found in Appendix A.1. Theorem 4.2 implies the following corollary for Gaussian likelihoods.
Corollary 4.1.
If $K=2$ and the likelihood is Gaussian, it holds that
$\lambda_{m_{1}}^{\text{SSOR}}=\dots=\lambda_{1}^{\text{SSOR}}=\lambda_{\max}^{\text{SSOR}}=1,$
$1-\left(\frac{\max(\sigma^{2}_{1},\sigma^{2}_{2})d^{\max}}{\sigma^{2}+\max(\sigma^{2}_{1},\sigma^{2}_{2})d^{max}}\right)^{2}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\left(\frac{\min(\sigma^{2}_{1},\sigma^{2}_{2})d^{\min}}{\sigma^{2}+\min(\sigma^{2}_{1},\sigma^{2}_{2})d^{\min}}\right)^{2},$ (25)
where $d^{\max}=\max(Z^{T}Z)$ and $d^{\min}=\min(Z^{T}Z)$ are the largest and smallest number of repeated measurements for the two grouped random effects, respectively.
If, in addition, the design of the random effects is balanced such that every random effect realization occurs $d$ times, i.e., $d=\frac{n}{m_{k}}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$ and $k=1,2$, the following holds true:
$\lambda_{\min}^{\text{SSOR}}$ $=1-\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{1}d}+1}\right)\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{2}d}+1}\right)$ (26)
Proof of Corollary 4.1.
The statement in (25) follows from (23). Similarly, (26) follows from (24) by plugging in $D_{k}=\text{diag}(1/\sigma^{2}_{k}+d_{k}/\sigma^{2})$, $k=1,2$. ∎
[…]
Theorem 4.3.
If $K=2$, the likelihood is Gaussian, and $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph with uniform distribution over all bipartite, biregular random graphs, the following hold:
$\lambda_{m-1}^{\text{SSOR}}\geq 1-\frac{(\sqrt{d_{1}-1}+\sqrt{d_{2}-1})^{2}}{(\sigma^{2}/\sigma_{1}^{2}+d_{1})(\sigma^{2}/\sigma_{2}^{2}+d_{2})}-\epsilon_{m}$ (27)
and
$\kappa_{m-1,1}^{\text{SSOR}}=\frac{\lambda_{1}^{\text{SSOR}}}{\lambda_{m-1}^{\text{SSOR}}}\leq\frac{1}{1-\left(\frac{1}{d_{1}}+\frac{1}{d_{2}}+\frac{2}{\sqrt{d_{1}d_{2}}}\right)-\epsilon_{m}}$ (28)
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ as $m\rightarrow\infty$, where $d_{k}=\frac{n}{m_{k}}$ for $k=1,2$.
A proof of Theorem 4.3 can be found in Appendix A.1. Pandolfi et al. (2024, Theorem 4) derive a similar result for the diagonal preconditioner. However, the upper bound for the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix given by $1/(1-(\frac{1}{d_{1}}+\frac{1}{d_{2}}+\frac{2}{\sqrt{d_{1}d_{2}}})-\epsilon_{m})$ decreases faster than the corresponding upper bound $(1+\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m})/(1-(\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}})-\epsilon_{m})$ for the diagonal preconditioner derived in Pandolfi et al. (2024, Theorem 4); see Section 4.3 for more details. Note that the random effects model used in Pandolfi et al. (2024) contains a random intercept term, which we do not include, and they derive their result for $\kappa_{m-2,3}^{\text{Diag}}$ instead of $\kappa_{m-1,1}^{\text{Diag}}$. However, in Theorem A.1 in Appendix A.2, we show that the same result holds true for $\kappa_{m-1,1}^{\text{Diag}}$ for the model without a random intercept used in this article.
4.2 Diagonal preconditioner and no preconditioner
We have the following results for the diagonal preconditioner and the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WT$. We comment on these results and compare them in Section 4.3.
Theorem 4.4.
For the diagonal preconditioner $P_{\text{Diag}}=D$, it holds that
$1+\frac{K-1}{\max(\Sigma_{ii}^{-1}(Z^{T}WZ)_{ii}^{-1})+1}\leq\lambda_{\max}^{\text{Diag}}\leq 1+\frac{K-1}{\min(\Sigma_{ii}^{-1}(Z^{T}WZ)_{ii}^{-1})+1},$ (29)
$\frac{1}{\max(\Sigma_{ii}D_{ii})}\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{\min(\Sigma_{ii}D_{ii})},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1.$ (30)
If the likelihood is Gaussian, we can additionally obtain the following result.
Theorem 4.5.
If the likelihood is Gaussian, it holds that
$1+\frac{K-1}{\frac{\max(\Sigma^{-1})\sigma^{2}}{d^{\min}}+1}\leq\lambda_{\max}^{\text{Diag}}\leq 1+\frac{K-1}{\frac{\min(\Sigma^{-1})\sigma^{2}}{d^{\max}}+1},$ (31)
$\frac{1}{\frac{\max(\Sigma)}{\sigma^{2}}d^{\max}+1}\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{\frac{\min(\Sigma)}{\sigma^{2}}d^{\min}+1},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1,$ (32)
where $d^{\max}=\max(Z^{T}Z)$ and $d^{\min}=\min(Z^{T}Z)$.
If, in addition, $K=2$ and the design of the random effects is balanced such that every random effect realization occurs $d_{k}$ times, i.e., $d_{k}=\frac{n}{m_{k}}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$ and $k=1,2$, it holds that
$\lambda_{\max}^{\text{Diag}}=1+\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1\right)}}$ (33)
$\lambda_{\min}^{\text{Diag}}=1-\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1\right)}}$ (34)
Theorem 4.6.
The following hold true for the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$:
$\frac{1}{\max(\Sigma_{ii})}+K\min((Z^{T}WZ)_{ii})\leq\lambda_{\max}^{\text{none}}\leq\frac{1}{\min(\Sigma_{ii})}+K\max((Z^{T}WZ)_{ii}),$ (35)
$\frac{1}{\max(\Sigma_{ii})}\leq\lambda_{m+1-k}^{\text{none}}\leq\frac{1}{\min(\Sigma_{ii})},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1.$ (36)
Proofs of Theorems 4.4, 4.5, and 4.6 can be found in Appendix A.1.
4.3 Comparison of preconditioners
The above results show that the largest and smallest eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for the SSOR, the diagonal, and no preconditioner depend only on the number of repeated observations per random effect, denoted by $d_{k}=\frac{n}{m_{k}}$ when having a balanced design, and not directly on neither the sample size $n$ nor the dimension $m$ of $\Sigma^{-1}+Z^{T}WZ$. In the following, we analyze and compare the extremal eigenvalues and condition numbers of the preconditioned matrices in more detail.
Theorem 4.7 below demonstrates that, for $K=2$, a Gaussian likelihood, and a balanced design with $d=m_{1}/m_{2}$ repeated observations per random effect, the condition numbers $\kappa=\lambda_{1}/\lambda_{m}$ for the SSOR and diagonal preconditioners and also when not applying any preconditioner grow asymptotically linearly in $d$. In detail, Theorem 4.7 shows that the condition number and the difference of the largest and smallest eigenvalue of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$, which matter for SLQ, are smaller for the SSOR preconditioner compared to the diagonal preconditioner. Furthermore, the theorem shows that the condition number $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ of the SSOR-preconditioned matrix has a linear asymptote in $d$ with a slope that is four times smaller than the corresponding slope of the diagonal preconditioner. In other words, the condition number of the diagonal-preconditioned matrix grows four times faster in $d$ than the one of the SSOR-preconditioned matrix. Based on this, we expect the SSOR preconditioner to result in more accurate SLQ-approximated log-determinants compared to the diagonal preconditioner. This is confirmed empirically in Section 5.
Furthermore, we can obtain an explicit expression for the condition number of the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ if the two random effects have equal variances, $\sigma^{2}_{1}=\sigma^{2}_{2}$ as the corresponding inequalities in Theorem 4.7 then become equalities. In this case, the condition number of the non-preconditioned and the diagonal-preconditioned matrix are equal and given by $\frac{2\sigma_{1}^{2}}{\sigma^{2}}d+1$, whereas the condition number of the SSOR-preconditioned matrix asymptotically grows with a rate of $\frac{\sigma^{2}_{1}}{2\sigma^{2}}$ in $d$. For the diagonal preconditioner, this can be seen by noting that $\lambda_{\max}^{\text{Diag}}/\lambda_{\min}^{\text{Diag}}=((\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1)+1)/((\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1)-1)=\frac{2\sigma_{1}^{2}}{\sigma^{2}}d+1$ due to (49) in the proof of Theorem 4.7 in Appendix A.1. We conjecture that similar results also hold for the non-preconditioned matrix and the diagonal preconditioner for the general case $\sigma^{2}_{1}\neq\sigma^{2}_{2}$.
Theorem 4.7.
If $K=2$, the likelihood is Gaussian, and the random effects have a balanced design with $d=d_{1}=\frac{n}{m_{1}}=d_{2}=\frac{n}{m_{2}}$ repeated observations per random effect, the following hold true:
$\lambda_{\max}^{\text{SSOR}}<\lambda_{\max}^{\text{Diag}},\leavevmode\nobreak\ \leavevmode\nobreak\ \lambda_{\min}^{\text{SSOR}}>\lambda_{\min}^{\text{Diag}},$ (37)
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}-\left(\frac{\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0,$ (38)
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}-\left(\frac{4\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0,$ (39)
$\frac{2\min(\sigma^{2}_{1},\sigma^{2}_{2})}{\sigma^{2}}d+\frac{\min(\sigma^{2}_{1},\sigma^{2}_{2})}{\max(\sigma^{2}_{1},\sigma^{2}_{2})}\leq\frac{\lambda_{\max}^{\text{none}}}{\lambda_{\min}^{\text{none}}}\leq\frac{2\max(\sigma^{2}_{1},\sigma^{2}_{2})}{\sigma^{2}}d+\frac{\max(\sigma^{2}_{1},\sigma^{2}_{2})}{\min(\sigma^{2}_{1},\sigma^{2}_{2})},$ (40)
We can additionally obtain upper bounds for the effective condition numbers $\kappa_{m-1,1}^{\text{SSOR}}$ and $\kappa_{m-1,2}^{\text{Diag}}$ for the SSOR and diagonal preconditioner, respectively, and a Gaussian likelihood when assuming that $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph. Theorem 4.8 below shows that the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix essentially decreases with a rate of $\frac{4}{d}$, up to higher-order terms. In contrast, the effective condition number $\kappa_{m-1,2}^{\text{Diag}}$ of the diagonal preconditioner decreases with a slower rate of $\frac{4}{\sqrt{d}}$, again up to higher-order terms. Note that the assumption $m_{1}=m_{2}$ in Theorem 4.8 allows for simpler results but can be easily relaxed. Based on these results, we expect the SSOR preconditioner to lead to faster convergence of the CG method compared to the diagonal preconditioner. A proof of Theorem 4.8 can be found in Appendix A.2.
Theorem 4.8.
If $K=2$, the likelihood is Gaussian, $m_{1}=m_{2}$, and $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph with uniform distribution over all bipartite, biregular random graphs, the following hold:
$\kappa_{m-1,1}^{\text{SSOR}}=\frac{\lambda_{1}^{\text{SSOR}}}{\lambda_{m-1}^{\text{SSOR}}}$ $\leq\frac{1}{1-\frac{4}{d}-\epsilon_{m}}=1+\frac{4}{d}+\epsilon_{m}+O\left(\frac{1}{d^{2}},\epsilon_{m}^{2},\frac{\epsilon_{m}}{d}\right),$ (41)
$\kappa_{m-1,2}^{\text{Diag}}=\frac{\lambda_{2}^{\text{Diag}}}{\lambda_{m-1}^{\text{Diag}}}$ $\leq\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}=1+\frac{4}{\sqrt{d}}+2\epsilon_{m}^{\prime}+O\left(\frac{1}{d},\epsilon_{m}^{\prime 2},\frac{\epsilon_{m}^{\prime}}{\sqrt{d}}\right)$ (42)
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ and $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$.
5 Experiments using simulated data
In the following, we conduct several experiments using simulated data. We compare the different preconditioners and methods for calculating predictive variances. Furthermore, we analyze our proposed methods concerning the accuracy and runtime for parameter estimation and prediction. Krylov subspace methods are compared to calculations based on the Cholesky decomposition using the GPBoost library. Note, however, that the Cholesky decomposition is also not exact since round-off errors can accumulate when using finite precision arithmetic. Additionally, we do a comparison with the lme4 and glmmTMB software libraries, which both rely on the Cholesky decomposition. Code to reproduce the simulated and the real-world data experiments of this article is available at https://github.com/pkuendig/KrylovGMMs.
5.1 Experimental setting
We simulate data from a model with two randomly crossed random effects. Unless stated otherwise, we use a balanced random effects design, in which both random effects have the same dimension $m_{1}=m_{2}$, and the random effect variances are $\sigma_{1}^{2}=\sigma_{2}^{2}=0.25$. But we also conduct experiments for an unbalanced design and other choices of random effects variances. We include a linear regression fixed effects term with five covariates plus an intercept for simulation, estimation, and prediction. The covariates are sampled from a normal distribution with mean zero and variance chosen such that the signal-to-noise ratio between the fixed and random effects is one, and the true regression coefficients are all one except for the intercept, which is zero. The response variable $y$ follows either a Gaussian likelihood with variance $\sigma^{2}=0.25$ or a Bernoulli likelihood with a logit link function. More details, e.g., on the sample size $n$ and the dimension $m$ of the random effects, are provided in the subsections below.
For analyzing prediction accuracy, the data is randomly split into training and test data sets. We measure both the accuracy of point and probabilistic predictions for the test data random effects $\tilde{b}_{p}=Z_{po}b+Z_{pp}b_{p}$; see (8). Point predictions in the form of predictive means are evaluated using the root mean squared error (RMSE), and for probabilistic predictions, we use the log score (LS) $-\frac{1}{n_{p}}\sum_{i=1}^{n_{p}}\log(\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii}))$, where $\omega_{p,i}$ and $(\Omega_{p})_{ii}$ are the predictive latent means and variances for the random effect $(\tilde{b}_{p})_{i}$ of test observation $i$, and $\mathcal{N}((\tilde{b}_{p})_{i};\omega_{p,i},(\Omega_{p})_{ii})$ denotes a Gaussian density evaluated at $(\tilde{b}_{p})_{i}$. We also measure the runtime in seconds.
All calculations are done on a laptop with an Intel i7-12800H processor and 32 GB of random-access memory. We use the GPBoost library version 1.5.8 for Krylov subspace and Cholesky decomposition-based methods denoted as ‘Krylov (GPBoost)’ and ‘Cholesky (GPBoost)’, respectively. Cholesky-based calculations are additionally done with the R packages lme4 version 1.1-35.5 and glmmTMB version 1.1.10. For Gaussian likelihoods, parameter estimation is done by minimizing the negative log-marginal likelihood. For non-Gaussian likelihoods, all implementations use the Laplace approximation. For parameter estimation, we set the initial values for all models to the default values used by GPBoost. Besides the above-mentioned points, the default options of the packages are used, including package-specific internal optimizers.
For the Krylov subspace methods, if not stated otherwise, we use the SSOR preconditioner, set the number of random vectors for STE and SLQ to $t=50$, and use a rank of $k=50$ for the two low-rank preconditioners. For the CG algorithm, we use a convergence tolerance of $10^{-2}$ for calculating marginal likelihoods and gradients, and a tolerance of $10^{-3}$ for predictive variances. Unless stated otherwise, we use the Algorithm 1 and $s=1000$ samples for calculating predictive variances. Further, we adopt a sample average approximation approach (Kim et al., 2015) when maximizing the marginal likelihood.
5.2 Preconditioner comparison
First, we compare the proposed preconditioners with regard to the accuracy and runtime for approximating log-marginal likelihoods. We set the total number of group levels to $m=100^{\prime}000$, use a sample size of $n=1^{\prime}000^{\prime}000$, and consider varying numbers of random vectors $t$ for the SLQ method. Note that these dimensions are approximately the largest ones such that Cholesky-based computations can be run in a “reasonable” amount of time which is roughly four hours for one likelihood evaluation. For each preconditioner and number of random vectors, the calculation of the likelihood is done at the true covariance parameters and repeated $100$ times with different random vectors. Figure 2 reports the results when the response variable follows a Gaussian likelihood. The runtime reported is the average wall-clock time. We first observe that the SSOR and ZIC preconditioners yield very accurate log-likelihood approximations with much smaller variances compared to the diagonal, Lanczos, and pivoted Cholesky preconditioners. This observation is consistent with the theoretical results presented in Section 4. Furthermore, we find that computations based on Krylov subspace methods are much faster than Cholesky-based computations. For instance, when using $t=50$ random vectors, we obtain a speed-up of approximately four orders of magnitude. Moreover, the runtimes of the CG method are similar for all preconditioners and even without preconditioning. In line with our convergence theory, the diagonal preconditioner has slightly longer runtimes compared to the SSOR preconditioner, but the differences are small. We conjecture that the finding that the preconditioned and non-preconditioned CG methods have similar runtimes can, partially, be explained by computational overheads of the preconditioners and other computations. The longest runtimes are observed for the Lanczos preconditioner. In summary, while preconditioning does not reduce the runtime, it can substantially reduce the variance of log-marginal likelihoods. In particular, variance reduction is very effective with the SSOR and ZIC preconditioner but has no clear effect with the diagonal, Lanczos, and pivoted Cholesky preconditioners.
Figure 2: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
In Figure 9 in Appendix A.8, we report analogous results when the response variable follows a Bernoulli likelihood with a logit link function and all other settings are identical for the SSOR, ZIC, and diagonal preconditioners. Similar observations can be made in this non-Gaussian setting, with the SSOR and ZIC preconditioners having substantially smaller variances than the diagonal preconditioner. Furthermore, in Figure 10 in Appendix A.8, we report the results when using an unbalanced random-effects design with $m_{2}=\frac{m_{1}}{2}$, and in Figure 11 in Appendix A.8 we present the results when having a different signal-to-noise ratio with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$ for a Gaussian likelihood. The results for these two settings are again very similar to the previous ones. Cholesky-based computations are faster for the unbalanced design, since $Z^{T}WZ$ is more sparse, but still three orders of magnitude slower than Krylov subspace methods. Note that the experiments in this section and the following Section 5.3 are conducted on only one simulated data set since we do not want to mix sampling variability and variability of the simulation-based approximations. However, the results do not change when using other samples (results not tabulated).
5.3 Comparing methods for predictive variances
Next, we compare the accuracy of the different approximations for predictive variances introduced in Section 3.2 using simulated data with $n=n_{p}=100^{\prime}000$ training and test points and a total of $m+m_{p}=10^{\prime}000$ group levels. The response variable follows a Gaussian likelihood, and we calculate the RMSE of the predictive variances compared to the “exact” Cholesky-based results. We also measure the runtime for prediction, which includes the calculation of the latent predictive means and variances. Predictive distributions are calculated at the true covariance parameters.
Figure 3 shows the RMSE versus the wall-clock time for different numbers of random vectors $s$ for the stochastic methods and different ranks $k$ for the Lanczos-based approximation. For the former, we average the results over $100$ independent repetitions for every $s$ and add whiskers to the plot representing confidence intervals for the RMSE obtained as ±2$\pm$ 2$\times$ 2 × standard errors, but the corresponding whiskers are not visible since they are very small. By far the most accurate predictive variances for a given runtime are obtained with the simulation-based Algorithm 1. For instance, for $s=200$, predictive variances are very accurate with an RMSE of approximately $2.6\times 10^{-4}$, and the runtime is more than $100$ times faster compared to Cholesky-based calculations. Algorithm 2 achieves the second highest prediction accuracy for a given runtime, and Algorithm 3 is considerably slower and less accurate. Finally, Lanczos-based approximations are very inaccurate and the RMSE decreases only very slowly with increasing rank $k$.
Figure 3: Comparison of different methods for predictive variances. The number of random vectors $s$ and the Lanczos rank $k$ are annotated in the plot.
5.4 Accuracy of parameter estimation and predictive distributions
In the following, we analyze the properties of the variance parameter and linear regression coefficient estimators and the accuracy of predictive distributions for Gaussian and Bernoulli likelihoods. We set the total number of group levels to $m+m_{p}=4^{\prime}000$ and simulate $n=n_{p}=40^{\prime}000$ training and test points. This is roughly the largest number of group levels such that we can run calculations with the glmmTMB and lme4 packages in a reasonable amount of time. We perform $100$ simulation repetitions, and prediction is done using estimated parameters. For Bernoulli likelihoods, repeating the estimation $100$ times with lme4 leads to very long runtimes; see also Section 5.5. Therefore, we do not report estimates and prediction accuracy for this case.
Figure 4 visualized the estimates for the variance parameters. The RMSE and bias of the variance and coefficient estimators are additionally reported in Appendix A.9 in Tables 2 and 3 for Gaussian and Bernoulli likelihoods, respectively. We observe virtually no differences among the estimates obtained using Krylov subspace methods and Cholesky-based computations with the GPBoost, lme4, and glmmTMB packages. Note that the variance parameter estimators for Bernoulli likelihoods are slightly downward biased due to the Laplace approximation and the moderate sample size.
Figure 4: Estimated variance parameters. The red rhombi represent means. The dashed lines indicate the true values. For the Bernoulli likelihood the estimates for lme4 are not computed due to excessively long runtimes.
Next, Figure 5 shows the RMSE for predictive means and the log score for predictive distributions. The average RMSE and the average log score with the corresponding standard errors are reported in Appendix A.9 in Tables 4 and 5 for Gaussian and Bernoulli likelihoods, respectively. Log scores are only reported for Krylov subspace methods and Cholesky-based computations from the GPBoost package since the calculation of predictive variances is not supported for lme4 and leads to very long runtimes for glmmTMB. We observe that the predictions obtained with the Cholesky decomposition and the Krylov subspace methods have virtually the same accuracy in terms of both the RMSE and the log score.
Figure 5: RMSE for predictive means and log score (LS) for probabilistic predictions for Gaussian and Bernoulli likelihoods.
5.5 Runtime comparison
In this subsection, we compare the runtime for parameter estimation when using Krylov subspace methods and Cholesky-based calculations with the GPBoost, lme4, and glmmTMB packages. For this, we consider different dimensions of the random effects ranging from $m=1^{\prime}000$ up to $m=1^{\prime}000^{\prime}000$ and $n=10m$ using the same setting as described in Section 5.1. Figure 6 shows the wall-clock time in seconds versus the total number of random effects $m$. Krylov subspace methods achieve the fastest runtime for all $m$. For instance, for $m=20^{\prime}000$ and a Gaussian likelihood, Krylov subspace methods are approximately three orders of magnitudes faster than Cholesky-based computations when using the GPBoost package. For a Gaussian likelihood, glmmTMB is approximately four orders of magnitudes slower than the Krylov subspace methods implemented in GPBoost, and lme4 is approximately four orders of magnitudes slower than the Krylov subspace methods for a Bernoulli likelihood. In general, estimation takes the longest with glmmTMB for Gaussian likelihoods, and for Bernoulli likelihoods, estimation takes the longest with lme4.
In Figure 12 in Appendix A.9, we additionally compare the runtime for parameter estimation for an unbalanced random effects design with $m_{2}=\frac{m_{1}}{2}$. Cholesky-based computations are faster compared to the balanced design due to the increased sparsity in $Z^{T}WZ$, but Krylov subspace methods again clearly achieve the fastest runtime for all $m$. For instance, they are three orders of magnitude faster than Cholesky-based computations using GPBoost for $m=50^{\prime}000$ and a Gaussian likelihood. Note that it would also be interesting to perform a runtime comparison for calculating log-marginal likelihoods as this would eliminate the impact of using different optimizers across different libraries, but the lme4 and glmmTMB packages do not support the explicit calculation of marginal likelihoods. However, ‘Krylov (GPBoost)’ and ‘Cholesky (GPBoost)’ differ only in the use of Krylov subspace methods vs. Cholesky-based computations, and this comparison therefore clearly demonstrates the speed-up with Krylov subspace methods.
Figure 6: Average wall clock times (s) for parameter estimation and different $m$. Simulated data follows either a Gaussian or a Bernoulli likelihood.
6 Real-world applications
In the following, we conduct experiments on five real-world data sets with high-dimensional crossed random effects analyzing the runtime and accuracy for parameter estimation when using GLMMs. We compare the proposed Krylov subspace methods with Cholesky-based calculations from the GPBoost, lme4, and glmmTMB packages, using the same model and computational settings as described in Section 5.1. We use the ZIC preconditioner for the regression data sets and the SSOR preconditioner for the classification data sets.
6.1 Data sets
We consider three regression and two classification data sets. Table 1 gives an overview of the data sets reporting the sample size, the number of predictor variables, the categorical variables modeled using random effects including the number of group levels $m_{j}$, and the total number of non-zero entries in $Z^{T}Z$. In Figure 13 in Appendix A.10, we additionally visualize the non-zero entries of the matrix $Z^{T}Z$ for the different data sets. Note that the black areas in the plots should not hide the fact that the high-dimensional matrices are sparse, which can be seen when the number of nonzero entries is set in relation to the matrix dimensions. The data sets ‘upselling’ (‘KDDCup09_upselling’) and ‘employee_access’ (‘Amazon_employee_access’) have previously been used by Pargent et al. (2022). For more information on the ‘instEval’ and ‘cars’ data sets, we refer to Simchoni and Rosset (2023), and the data set ‘building_permits’ (‘chicago_building_permits’) has been used by Reyes (2019). In general, an extra category is added for missing values in categorical variables and numeric predictor variables are simply imputed using the mean.
data set $n$ $p$ $K$ $y$-class Cat. var. $m_{j}$ $\text{nzz}(Z^{T}Z)$
Regression cars 97’729 66 2 log(price) model_id 15’226 185’807
location_id 12’235
building_permits 527’168 6 3 log(cost) contact_name 98’362 2’544’893
latitude 206’197
longitude 206’170
instEval 73’421 22 2 teacher student 2’972 150’942
rating teacher 1’128
Classification employee_access 32’769 0 9 approval 0.94 resource 7’518 469’262
manager 4’243
role_cat1 128
role_cat2 177
role_dep. 449
role_title 343
role_descr. 2’358
role_family 67
role_code 343
upselling 50’000 34 4 up-selling 0.07 Var216 2’016 454’852
Var217 13’991
Var198 4’291
Var199 5’074
Table 1: Summary of real-world data sets. $n$ is the number of samples, $p$ is the number of predictor variables, $K$ is the number of categorical variables modeled with random effects, $y$ describes the response variable, ‘%-class’ is the frequency of the ‘1’ response variables for the classification data sets, ‘Cat. var.’ describes the categorical variables modeled with random effects, $m_{j}$ is the number of group levels, and $\text{nzz}(Z^{T}Z)$ is the number of non-zero entries in the matrix $Z^{T}Z$.
6.2 Results
Figure 7 shows the estimation runtimes for the different data sets. In Appendix A.10, we additionally report the wall-clock time, the log-marginal likelihood at the optimum, and the estimated parameters. Parameter estimation crashes when using the lme4 and glmmTMB packages on the ‘upselling’ data set. We have additionally tried modeling only the two variables Var216 and Var217 as random effects but still observed crashes on the ‘upselling’ data set. For the ‘employee_access’ data set, we have also tried modeling the low-cardinality categorical variables such as ‘role_family’ and ‘role_cat1’ as dummy-coded fixed effects instead of random effects, but glmmTMB and lme4 crash when doing this. In Figure 8, we report the average differences in estimation runtime relative to the fastest model over the four data sets for which glmmTMB and lme4 do not crash.
Figure 7: Runtimes for estimation on different real-world data sets for different models.
The results show that Krylov subspace methods clearly have the fastest runtime on all real-world data sets. Compared to Cholesky-based calculations using the GPBoost package, the average speed-up with Krylov subspace methods is more than a factor of $20$. On the high-dimensional ‘building_permits’ data set, parameter estimation with Krylov subspace methods is over 57 times faster compared to Cholesky-based calculations in GPBoost. Moreover, estimation is, on average, more than $30$ and $80$ times slower when using lme4 and glmmTMB, respectively, compared to Krylov subspace-based methods. On average, the estimation time of lme4 is the second smallest for the three regression data sets. However, as observed in experiments with simulated data, lme4 is very slow on classification data and can crash, e.g., on the ‘upselling’ data set. glmmTMB has overall the longest runtime. Concerning the parameter estimates reported in Tables 8 and 7 in Appendix A.10, we find that all methods and packages yield essentially the same estimates except for lme4 on the ‘building_permits’ and ‘employee_access’ data sets. In summary, our novel Krylov subspace-based methods are substantially faster than Cholesky-based calculations and numerically more stable than existing software libraries for GLMMs.
Figure 8: Average relative difference in runtime to the fastest model over all real-world data sets except for the data set ‘upselling’ where lme4 and glmmTMB crash.
7 Conclusion
We present Krylov subspace methods for mixed effects models with high-dimensional crossed random effects. We analyze several preconditioners, derive novel theoretical results, and introduce different methods for accurately approximating predictive variances. In experiments, we find that our methods are faster and more stable than Cholesky-based calculations and other state-of-the-art software implementations for GLMMs. In addition, we provide theoretical and empirical results showing that the symmetric successive over-relaxation (SSOR) preconditioner is superior compared to the popular diagonal preconditioner for both the CG and the SLQ methods. Potential directions for future research include analyzing alternative approximations for log-determinants, a systematic comparison under which situations the ZIC or the SSOR preconditioners are preferred, and the development of novel preconditioners.
Acknowledgments
This research was partially supported by the Swiss Innovation Agency - Innosuisse (grant number ‘55463.1 IP-ICT’).
Appendix A Appendix
A.1 Proofs of results in Section 4
We denote by $\sigma_{\min}(B)=\sigma_{\min(n_{1},n_{2})}(B)\leq\dots\leq\sigma_{1}(B)=\sigma_{\max}(B)$ the singular values of a matrix $B\in\mathbb{R}^{n_{1}\times n_{2}}$, and by $\lambda_{\min}(A)=\lambda_{m}(A)\leq\dots\leq\lambda_{1}(A)=\lambda_{\max}(A)$ the eigenvalues of a symmetric matrix $A\in\mathbb{R}^{m\times m}$. Further, we denote by $D=\Sigma^{-1}+\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$, $\Delta=\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$, and $\Lambda=Z^{T}WZ-\text{diag}(Z^{T}WZ)\in\mathbb{R}^{m\times m}$.
We will use the following two lemmas for proofing Theorem 4.2.
Lemma A.1.
For any $B\in\mathbb{R}^{n_{1}\times n_{2}}$ and $A=\begin{pmatrix}0&B^{T}\\ B&0\end{pmatrix}\in\mathbb{R}^{(n_{1}+n_{2})\times(n_{1}+n_{2})}$, the non-zero eigenvalues of $A$ are equal to $\pm$ the singular values of $B$, i.e.,
$\{\lambda_{i}(A);i=1,\dots,2\min(n_{1},n_{2})\}=\{\pm\sigma_{i}(B);i=1,\dots,\min(n_{1},n_{2})\}.$
In particular, the largest singular value of $B$ and the largest eigenvalue of $A$ are equal, $\lambda_{\max}(A)=\sigma_{\max}(B)$. The same holds true for the second largest singular value and eigenvalue.
Proof of Lemma A.1.
The eigenvalues of $A$ are defined as solutions to
$\text{det}(A-\lambda I)=0.$
Since $A-\lambda I$ is a block matrix, we have
$\text{det}(A-\lambda I)=\text{det}(\lambda^{2}I-B^{T}B).$
The eigenvalues of $A$ are thus given by
{±$\{\pm\sqrt{\lambda_{i}(B^{T}B)};i=1,\dots,\min(n_{1},n_{2})\}=\{\pm\sigma_{i}(B);i=1,\dots,\min(n_{1},n_{2})\}.$
Lemma A.2.
For any matrix $B\in\mathbb{R}^{n_{1}\times n_{2}}$ and diagonal matrices $C_{1}^{n_{1}\times n_{1}}$ and $C_{2}^{n_{2}\times n_{2}}$, it holds that
$\sigma_{\max}(B)\min(|(C_{1})_{ii}|)\min(|(C_{2})_{jj}|)\leq\sigma_{\max}(C_{1}BC_{2})\leq\sigma_{\max}(B)\max(|(C_{1})_{ii}|)\max(|(C_{2})_{jj}|).$
Proof of Lemma A.2.
$\sigma_{\max}(C_{1}BC_{2})=\|C_{1}BC_{2}\|_{2}\leq\|C_{1}\|_{2}\|B\|_{2}\|C_{2}\|_{2}=\sigma_{\max}(B)\max(|(C_{1})_{ii}|)\max(|(C_{2})_{jj}|).$
Furthermore,
$\sigma_{\max}(B)=\|B\|_{2}=\|C_{1}^{-1}C_{1}BC_{2}C_{2}^{-1}\|_{2}\leq\sigma_{\max}(C_{1}BC_{2})\frac{1}{\min(|(C_{1})_{ii}|)}\frac{1}{\min(|(C_{2})_{ii}|)}$
which gives the first inequality. ∎
Proof of Theorem 4.2.
For the SSOR preconditioner $P_{\text{SSOR}}=(L+D)D^{-1}(L+D)^{T}$, we have
$L=\begin{pmatrix}0&0\\ Z_{2}^{T}WZ_{1}&0\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ D=\begin{pmatrix}D_{1}&0\\ 0&D_{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \Sigma^{-1}+Z^{T}WZ=\begin{pmatrix}D_{1}&Z_{1}^{T}WZ_{2}\\ Z_{2}^{T}WZ_{1}&D_{2}\end{pmatrix}.$
It can easily be seen that
$(D+L)^{-1}=\begin{pmatrix}D_{1}^{-1}&0\\ -D_{2}^{-1}Z_{2}^{T}WZ_{1}D_{1}^{-1}&D_{2}^{-1}\end{pmatrix},$
and (21) thus follows since
$\begin{split}P_{\text{SSOR}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{SSOR}}^{-T/2}&=D^{1/2}(D+L)^{-1}(\Sigma^{-1}+Z^{T}WZ)(D+L)^{-T}D^{1/2}\\ &=I_{m}-\begin{pmatrix}0&0\\ 0&D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}\end{pmatrix}.\end{split}$ (43)
Because $D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}$ is a symmetric positive semi-definite matrix and using the fact that $\lambda_{i}(I_{m}+A)=1+\lambda_{i}(A)$ for any normal matrix $A$, we obtain (22).
Using Lemma A.1, we have
$\begin{split}\sigma_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1/2})&=\lambda_{\max}\begin{pmatrix}0&D_{1}^{-1/2}Z_{1}^{T}WZ_{2}D_{2}^{-1/2}\\ D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1/2}&0\end{pmatrix}\\ &=\lambda_{\max}(D^{-1/2}\Lambda D^{-1/2})\\ &=\lambda_{\max}(D^{-1/2}\Delta^{1/2}\Delta^{-1/2}\Lambda\Delta^{-1/2}\Delta^{1/2}D^{-1/2}).\end{split}$
Note that $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is a normalized adjacency matrix of a bipartite weighted graph. Its largest eigenvalues is thus $1$ [Chung, 1997]. By Lemma A.2, or Ostrowski’s theorem [Horn and Johnson, 2012], it follows that
$\min((D^{-1}\Delta)_{ii})\leq\sigma_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1/2})\leq\max((D^{-1}\Delta)_{ii}).$
Since $D^{-1}\Delta=\text{diag}\left(\frac{1}{\Sigma^{-1}_{ii}\Delta^{-1}_{ii}+1}\right)$ and
[…]
we have
$\frac{1}{(\max(\Sigma^{-1}_{ii}\Delta^{-1}_{ii})+1)^{2}}\leq\lambda_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2})\leq\frac{1}{(\min(\Sigma^{-1}_{ii}\Delta^{-1}_{ii})+1)^{2}},$
from which follows (23).
Furthermore, since
$\begin{split}\sigma_{\max}(Z_{2}^{T}WZ_{1})&=\lambda_{\max}\begin{pmatrix}0&Z_{1}^{T}WZ_{2}\\ Z_{2}^{T}WZ_{1}&0\end{pmatrix}=\lambda_{\max}(\Lambda)=\lambda_{\max}(\Delta^{1/2}\Delta^{-1/2}\Lambda\Delta^{-1/2}\Delta^{1/2}).\end{split}$
and applying Lemma A.2, we obtain
$\min(\Delta_{ii})\leq\sigma_{\max}(Z_{2}^{T}WZ_{1})\leq\max(\Delta_{ii}).$
By again using Lemma A.2, we obtain
$\min((D_{1}^{-1/2})_{ii})\min((D_{2}^{-1/2})_{ii})\min(\Delta_{ii})\leq\sigma_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1/2})\leq\max((D_{1}^{-1/2})_{ii})\max((D_{2}^{-1/2})_{ii})\max(\Delta_{ii})$
and
$\min((D_{1}^{-1})_{ii})\min((D_{2}^{-1})_{ii})\min(\Delta_{ii}^{2})\leq\lambda_{\max}(D_{2}^{-1/2}Z_{2}^{T}WZ_{1}D_{1}^{-1}Z_{1}^{T}WZ_{2}D_{2}^{-1/2})\leq\max((D_{1}^{-1})_{ii})\max((D_{2}^{-1})_{ii})\max(\Delta_{ii}^{2})$
which gives the second set of inequalities for $\lambda_{\min}^{\text{SSOR}}$ in (24). ∎
Proof of Theorem 4.3.
Similarly as in the proof of Theorem 4.2, we can apply Lemma A.1 and obtain
$\begin{split}\sigma_{2}\left(\frac{1}{\sigma^{2}}D_{2}^{-1/2}Z_{2}^{T}Z_{1}D_{1}^{-1/2}\right)&=\lambda_{2}\left(\frac{1}{\sigma^{2}}D^{-1/2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}D^{-1/2}\right)\\ &=\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\frac{1}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}}\frac{1}{\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}},\end{split}$ (45)
where we have used
$\sigma^{-1}D^{-1/2}=\text{diag}\left(\frac{1}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}},\dots,\frac{1}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}},\frac{1}{\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}},\dots,\frac{1}{\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}}\right).$
Applying Theorem 3.2 of Brito et al. [2022] to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
$\begin{split}\sigma_{2}\left(\frac{1}{\sigma^{2}}D_{2}^{-1/2}Z_{2}^{T}Z_{1}D_{1}^{-1/2}\right)&\leq\frac{\sqrt{d_{1}-1}+\sqrt{d_{2}-1}}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}}+\epsilon_{m}^{\prime}\\ &\leq\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m}^{\prime}\end{split}$
asymptotically almost surely with $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$. Using (43) and (44) proofs (27). (28) follows directly from (27).
Proof of Theorem 4.4.
Consider the vectors
$v_{k}=D^{1/2}(\frac{1}{k}1_{m_{1}}^{T},\dots,\frac{1}{k}1_{m_{k}}^{T},-1_{m_{k+1}}^{T},0,\dots,0)^{T}\in\mathbb{R}^{m},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1,$
where $1_{m_{k}}$ denotes a vectors of $1$’s of length $m_{k}$. We have $D^{-1/2}Z^{T}WZD^{-1/2}v_{k}=0$, and the vectors $v_{k}$ are linearly independent. Since $D^{-1/2}Z^{T}WZD^{-1/2}$ is positive semi-definite, this shows that the $K-1$ smallest eigenvalues of $D^{-1/2}Z^{T}WZD^{-1/2}$ are zero. Applying Weyl’s inequality [Horn and Johnson, 2012] to $P_{\text{Diag}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{Diag}}^{-T/2}=D^{-1/2}\Sigma^{-1}D^{-1/2}+D^{-1/2}Z^{T}WZD^{-1/2}$ gives
$\lambda_{\min}(D^{-1/2}\Sigma^{-1}D^{-1/2})\leq\lambda_{m+1-k}(P_{\text{Diag}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{Diag}}^{-T/2})\leq\lambda_{\max}(D^{-1/2}\Sigma^{-1}D^{-1/2})$
for $k=1,\dots,K-1$, which proofs (30).
Similarly as in the proof of Theorem 4.2, we can show, for general $K$, that
$(K-1)\min((D^{-1}\Delta)_{ii})\leq\lambda_{\max}(D^{-1/2}\Lambda D^{-1/2})\leq(K-1)\max((D^{-1}\Delta)_{ii}).$ (46)
For doing this, we first note that $\Lambda$ is an adjacency matrix of a connected, $K$-partite weighted graph. Moreover,
$\begin{split}\sum_{j=1}^{m}\Lambda_{ij}&=\sum_{k=1,k\neq\tilde{k}(i)}^{K}\sum_{j=1}^{m_{k}}(Z_{\tilde{k}(i)}^{T}WZ_{k})_{ij}\\ &=\sum_{k=1,k\neq\tilde{k}(i)}^{K}\sum_{j=1}^{m_{k}}\sum_{l=1}^{n}W_{ll}(Z_{\tilde{k}(i)})_{li}(Z_{k})_{lj}\\ &=\sum_{k=1,k\neq\tilde{k}(i)}^{K}\sum_{l=1}^{n}W_{ll}(Z_{\tilde{k}(i)})_{li}\\ &=\sum_{k=1,k\neq\tilde{k}(i)}^{K}\Delta_{ii}=(K-1)\Delta_{ii}\end{split}$
where $\tilde{k}(i)=k$ if $i\in\{\sum_{k^{\prime}=0}^{k-1}m_{k^{\prime}}+1,\dots,\sum_{k^{\prime}=0}^{k}m_{k^{\prime}}\}$ with the convention $m_{0}=0$. This shows that $\frac{1}{K-1}\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is a normalized adjacency matrix of a connected, $K$-partite weighted graph whose largest eigenvalue is $1$. Using
$\lambda_{\max}(D^{-1/2}\Lambda D^{-1/2})=(K-1)\lambda_{\max}(D^{-1/2}\Delta^{1/2}\frac{1}{K-1}\Delta^{-1/2}\Lambda\Delta^{-1/2}\Delta^{1/2}D^{-1/2})$
gives (46), which together with
$P_{\text{Diag}}^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P_{\text{Diag}}^{-T/2}=I_{m}+D^{-1/2}\Lambda D^{-1/2}$ (47)
proofs the statement in (29). ∎
Proof of Theorem 4.5.
First, (31) and (32) follow directly from (29) and (30).
For (33), note that
$\begin{split}D^{-1/2}\Delta^{1/2}&=\text{diag}\left(\frac{1}{\Sigma^{-1}_{ii}\Delta^{-1}_{ii}+1}\right)^{1/2}\\ &=\text{diag}\left(\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1}},\dots,\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1}},\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1}},\dots,\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1}}\right),\end{split}$
which implies
$D^{-1/2}\Delta^{1/2}\Delta^{-1/2}\Lambda\Delta^{-1/2}\Delta^{1/2}D^{-1/2}=\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1}}\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1}}\Delta^{-1/2}\Lambda\Delta^{-1/2}.$ (48)
Using $\lambda_{\max}(\Delta^{-1/2}\Lambda\Delta^{-1/2})=K-1$, see the proof of Theorem 4.4, then gives
$\begin{split}\lambda_{\max}(D^{-1/2}\Lambda D^{-1/2})&=\lambda_{\max}(D^{-1/2}\Delta^{1/2}\Delta^{-1/2}\Lambda\Delta^{-1/2}\Delta^{1/2}D^{-1/2})\\ &=\frac{K-1}{\sqrt{\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1}\sqrt{\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1}}.\end{split}$
This and (47) proof (33).
By analogous arguments, again using (48) and $\lambda_{\min}(\Delta^{-1/2}\Lambda\Delta^{-1/2})=-1$ due to Lemma A.1, we have
$\lambda_{\min}(D^{-1/2}\Lambda D^{-1/2})=-\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{1}^{2}d_{1}}+1}}\frac{1}{\sqrt{\frac{\sigma^{2}}{\sigma_{2}^{2}d_{2}}+1}},$
which gives (34). ∎
Proof of Theorem 4.6.
We have
$\begin{split}\Sigma^{-1}+Z^{T}WZ=\Sigma^{-1}+\Delta^{1/2}(I_{m}+\Delta^{-1/2}\Lambda\Delta^{-1/2})\Delta^{1/2}.\end{split}$
Applying Weyl’s inequality [Horn and Johnson, 2012], Ostrowski’s theorem [Horn and Johnson, 2012], and the fact that the largest eigenvalue of $\Delta^{-1/2}\Lambda\Delta^{-1/2}$ is $K-1$, see the proof of Theorem 4.4, we obtain (35). (36) is shown analogously as (30) in the proof of Theorem 4.4. ∎
Proof of Theorem 4.7.
First, (37) follows directly from Corollary 4.1 and Theorem 4.5. Next, by Corollary 4.1, we have
$\lambda_{\min}^{\text{SSOR}}=1-\frac{\sigma^{2}_{1}d}{\sigma^{2}+\sigma^{2}_{1}d}\frac{\sigma^{2}_{2}d}{\sigma^{2}+\sigma^{2}_{2}d}=\frac{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d}{(\sigma^{2})^{2}+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})d+\sigma^{2}_{1}\sigma^{2}_{2}d^{2}}$
and thus
$\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}=\frac{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})+\sigma^{2}_{1}\sigma^{2}_{2}d}{(\sigma^{2})^{2}/d+\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})},$
and (38) follows. By Theorem 4.5, we have
$\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}=\frac{\left(1+\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}{\left(1-\frac{1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}}\right)}=\frac{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}+1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}-1}.$ (49)
Using
$\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}=1+\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+\frac{\sigma^{2}}{\sigma_{2}^{2}d}\right)+\mathcal{O}(d^{-2}),$
one can show that $\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}$ has the linear asymptote in $d$ given in (39). (40) follows from Theorem 4.6.
A.2 Effective condition number for the diagonal preconditioner and proof of Theorem 4.8
Theorem A.1.
If $K=2$, the likelihood is Gaussian, and $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph with uniform distribution over all bipartite, biregular random graphs, the following holds:
$\kappa_{m-1,2}^{\text{Diag}}=\frac{\lambda_{2}^{\text{Diag}}}{\lambda_{m-1}^{\text{Diag}}}\leq\frac{1+\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m}}{1-\left(\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}\right)-\epsilon_{m}}$ (50)
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ as $m\rightarrow\infty$, where $d_{k}=\frac{n}{m_{k}}$ for $k=1,2$.
Proof of Theorem A.1.
We have $\lambda_{2}^{\text{Diag}}=1+\lambda_{2}(D^{-1/2}\Lambda D^{-1/2})$ because of (47). Next, (45) gives
$\lambda_{2}(D^{-1/2}\Lambda D^{-1/2})=\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\frac{1}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}}\frac{1}{\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}}.$
Analogously as in the proof of Theorem 4.3, applying Theorem 3.2 of Brito et al. [2022] to $\lambda_{2}\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}$ then shows that
$\begin{split}\lambda_{2}^{\text{Diag}}&\leq 1+\frac{\sqrt{d_{1}-1}+\sqrt{d_{2}-1}}{\sqrt{\sigma^{2}/\sigma_{1}^{2}+d_{1}}\sqrt{\sigma^{2}/\sigma_{2}^{2}+d_{2}}}+\epsilon_{m}\\ &\leq 1+\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m}\end{split}$
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ as $m\rightarrow\infty$. Lemma A.1 and analogous arguments give
$\lambda_{m-1}^{\text{Diag}}\geq 1-\left(\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}\right)-\epsilon_{m}.$
The above two inequalities then result in (50). ∎
Proof of Theorem 4.8.
Theorem 4.3 shows that
$\kappa_{m-1,1}^{\text{SSOR}}\leq\frac{1}{1-\frac{4}{d}-\epsilon_{m}}.$
Applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ with $x=\frac{4}{d}+\epsilon_{m}$ shows (41). Similarly, Theorem A.1, applying the expansion $(1-x)^{-1}=1+x+x^{2}+O(x^{3})$ in the denominator of $\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}$, and multiplying by the numerator $1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}$ give (42). ∎
A.3 Proofs of Proposition 3.1, Proposition 3.2, and Proposition 3.3
Proof of Proposition 3.1.
By standard results, $\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(1)}\odot z_{i}^{(2)}=\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(1)}\odot Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}z_{i}^{(1)}$ in Algorithm 1 is an unbiased and consistent estimator for $\text{diag}(Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})$ and $\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(1)}\odot z_{i}^{(3)}=\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(1)}\odot Z_{po}P^{-1}Z_{po}^{T}z_{i}^{(1)}$ is an unbiased and consistent estimator for $\text{diag}(Z_{po}P^{-1}Z_{po}^{T})$. Thus, the claim in Proposition 3.1 follows. ∎
Proof of Proposition 3.2.
First, observe that $z_{i}^{(3)}=\Sigma^{-\frac{1}{2}}z_{i}^{(1)}+Z^{T}W^{\frac{1}{2}}z_{i}^{(2)}\sim\mathcal{N}(0,(\Sigma^{-1}+Z^{T}WZ))$. It follows that $z_{i}^{(4)}=Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}^{(3)}\sim\mathcal{N}(0,Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T})$. By standard results, $\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$ in Algorithm 2 is an unbiased and consistent estimator for $Z_{po}(\Sigma^{-1}+Z^{T}WZ)^{-1}Z_{po}^{T}$, and the claim in Proposition 3.2 thus follows. ∎
Proof of Proposition 3.3.
First, observe that $z_{i}^{(3)}=Z\Sigma^{\frac{1}{2}}z_{i}^{(1)}+W^{-\frac{1}{2}}z_{i}^{(2)}\sim\mathcal{N}(0,\Psi)$. It follows that $z_{i}^{(4)}=Z_{po}\Sigma Z^{T}\Psi^{-1}z_{i}^{(3)}\sim\mathcal{N}(0,Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T})$. By standard results, $\frac{1}{s}\sum_{i=1}^{s}z_{i}^{(4)}\left(z_{i}^{(4)}\right)^{T}$ in Algorithm 3 is an unbiased and consistent estimator for $Z_{po}\Sigma Z^{T}\Psi^{-1}Z\Sigma Z_{po}^{T}$, and the claim in Proposition 3.3 thus follows. ∎
A.4 Derivatives of log-determinants using stochastic trace estimation and variance reduction for SSOR preconditioner
In the following, $c=\widehat{\text{Cov}}(h(z_{i}),r(z_{i}))/\widehat{\text{Var}}(r(z_{i}))$ is the optimal weight for the variance reduction.
A.4.1 Derivative for variance parameters
Note that $\frac{\partial D}{\partial\theta_{k}}=\frac{\partial\Sigma^{-1}}{\partial\theta_{k}}=-\Sigma^{-1}\Sigma^{-1}$.
$\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}\approx c\;\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\theta_{k}}}_{\text{deterministic}}+\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}-c\;\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\theta_{k}}}_{\text{stochastic}}$
$\begin{split}\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\theta_{k}}}_{\text{deterministic}}&=\frac{\partial\left(\log\det(L+D)+\log\det(D^{-1})+\log\det(L+D)\right)}{\partial\theta_{k}}\\ &=\frac{\partial\log\det(D)}{\partial\theta_{k}}\\ &=\operatorname*{tr}\left(D^{-1}\frac{\partial D}{\partial\theta_{k}}\right)\end{split}$
$\begin{split}\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}=&\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\theta_{k}}\right)\\ =&\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial\Sigma^{-1}}{\partial\theta_{k}}\right)\\ =&\operatorname*{tr}\left(-(\Sigma^{-1}+Z^{T}WZ)^{-1}\Sigma^{-1}\Sigma^{-1}\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\underbrace{-\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}\Sigma^{-1}\Sigma^{-1}P_{\text{SSOR}})^{-1}z_{i}}_{=:h(z_{i})}\end{split}$
$\begin{split}\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\theta_{k}}}_{\text{stochastic}}=&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\frac{\partial P_{\text{SSOR}}}{\partial\theta_{k}}\right)\\ =&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\frac{\partial(L+D)D^{-1}(L+D)^{T}}{\partial\theta_{k}}\right)\\ =&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\left(2\;\frac{\partial D}{\partial\theta_{k}}D^{-1}(L+D)^{T}-(L+D)D^{-1}\frac{\partial D}{\partial\theta_{k}}D^{-1}(L+D)^{T}\right)\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\underbrace{\left(P_{\text{SSOR}}^{-1}z_{i}\right)^{T}\left(2\;\frac{\partial D}{\partial\theta_{k}}D^{-1}(L+D)^{T}-(L+D)D^{-1}\frac{\partial D}{\partial\theta_{k}}D^{-1}(L+D)^{T}\right)P_{\text{SSOR}}^{-1}z_{i}}_{=:r(z_{i})}\end{split}$
A.4.2 Further derivatives for Laplace approximations
For efficient calculations of the derivatives of log-determinants with STE, note that $\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial F_{i}}=\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}$ and $\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial b^{*}_{j}}=\left(\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}}\right)^{T}\frac{\partial\mu^{*}}{\partial b^{*}_{j}}=\left(\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}}\right)^{T}Z_{j}$. Therefore, we present in the following calculations for $\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}$, and $\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\xi_{l}}$ is obtained by replacing $\frac{\partial W}{\partial\mu^{*}_{i}}=\text{diag}(-\frac{\partial^{3}\log p(y_{i}|\mu^{*}_{i},\xi)}{\partial{\mu^{*}_{i}}^{3}})$ with $\frac{\partial W}{\partial\xi_{l}}=\text{diag}(-\frac{\partial^{3}\log p(y_{i}|\mu^{*}_{i},\xi)}{\partial{\mu^{*}_{i}}^{2}\partial\xi_{l}})$. Further, $\frac{\partial D}{\partial\mu^{*}_{i}}$ is a diagonal matrix with diagonal entries $\left(\frac{\partial D}{\partial\mu^{*}_{i}}\right)_{ii}=\left(Z^{T}\frac{\partial W}{\partial\mu^{*}_{i}}Z\right)_{ii}$, and $\frac{\partial(L+D)}{\partial\mu^{*}_{i}}$ is a lower-triangular matrix with entries $\left(\frac{\partial(L+D)}{\partial\mu^{*}_{i}}\right)_{ij}=\mathbf{1}_{\{i=>j\}}\left(Z^{T}\frac{\partial W}{\partial\mu^{*}_{i}}Z\right)_{ij}$.
$\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}\approx c\;\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\mu^{*}_{i}}}_{\text{deterministic}}+\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}-c\;\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\mu^{*}_{i}}}_{\text{stochastic}}$
$\begin{split}\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\mu^{*}_{i}}}_{\text{deterministic}}=&\frac{\partial\left(\log\det(L+D)+\log\det(D^{-1})+\log\det(L+D)\right)}{\partial\mu^{*}_{i}}\\ =&\frac{\partial\log\det(D)}{\partial\mu^{*}_{i}}\\ =&\operatorname*{tr}\left(D^{-1}\frac{\partial D}{\partial\mu^{*}_{i}}\right)\end{split}$
$\begin{split}\frac{\partial\log\det(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}=&\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}\frac{\partial(\Sigma^{-1}+Z^{T}WZ)}{\partial\mu^{*}_{i}}\right)\\ =&\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}\frac{\partial W}{\partial\mu^{*}_{i}}Z\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\underbrace{\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}\frac{\partial W}{\partial\mu^{*}_{i}}ZP_{\text{SSOR}}^{-1}z_{i}}_{=:h(z_{i})}\end{split}$
$\begin{split}\underbrace{\frac{\partial\log\det(P_{\text{SSOR}})}{\partial\mu^{*}_{i}}}_{\text{stochastic}}=&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\frac{\partial P_{\text{SSOR}}}{\partial\mu^{*}_{i}}\right)\\ =&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\frac{\partial(L+D)D^{-1}(L+D)^{T}}{\partial\mu^{*}_{i}}\right)\\ =&\operatorname*{tr}\left(P_{\text{SSOR}}^{-1}\left(2\;\frac{\partial(L+D)}{\partial\mu^{*}_{i}}D^{-1}(L+D)^{T}-(L+D)D^{-1}\frac{\partial D}{\partial\mu^{*}_{i}}D^{-1}(L+D)^{T}\right)\right)\\ \approx&\frac{1}{t}\sum_{i=1}^{t}\underbrace{\left(P_{\text{SSOR}}^{-1}z_{i}\right)^{T}\left(2\;\frac{\partial(L+D)}{\partial\mu^{*}_{i}}D^{-1}(L+D)^{T}-(L+D)D^{-1}\frac{\partial D}{\partial\mu^{*}_{i}}D^{-1}(L+D)^{T}\right)P_{\text{SSOR}}^{-1}z_{i}}_{=:r(z_{i})}\end{split}$
A.5 Alternative stochastic trace estimation for Fisher information
For $1\leq k,l\leq K$ and $z_{i}\sim\mathcal{N}(0,P)$, the trace terms of the Fisher information in (7) can alternatively be computed with STE as follows:
$\begin{split}&\operatorname*{tr}\left(\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{k}}\Psi^{-1}\frac{\partial\Psi}{\partial\theta_{l}}\right)\\ =&\operatorname*{tr}\left(\left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{k}Z_{k}^{T}\right.\\ &\left.\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \left(W-WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}W\right)Z_{l}Z_{l}^{T}\right)\\ =&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ &+\leavevmode\nobreak\ \leavevmode\nobreak\ \operatorname*{tr}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZW\right)\\ \approx&\operatorname*{tr}\left(WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}\right)\\ &-2\frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}\\ &+\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{1}{t}\sum_{i=1}^{t}\left((\Sigma^{-1}+Z^{T}WZ)^{-1}z_{i}\right)^{T}Z^{T}WZ_{k}Z_{k}^{T}WZ(\Sigma^{-1}+Z^{T}WZ)^{-1}Z^{T}WZ_{l}Z_{l}^{T}ZWP^{-1}z_{i}.\end{split}$ (51)
A.6 Preconditioned conjugate gradient algorithm
Algorithm 4 Preconditioned conjugate gradient algorithm with Lanczos tridiagonal matrix
1:Matrix $A$, preconditioner matrix $P$, vector $b$
2:$u_{l+1}\approx A^{-1}b$, tridiagonal matrix $\tilde{T}$
3:early-stopping $\leftarrow$ false
4:$\alpha_{0}\leftarrow 1$
5:$\beta_{0}\leftarrow 0$
6:$u_{0}\leftarrow 0$
7:$r_{0}\leftarrow b-Au_{0}$
8:$z_{0}\leftarrow P^{-1}r_{0}$
9:$h_{0}\leftarrow z_{0}$
10:for $l\leftarrow 0$ to $L$ do
11:     $v_{l}\leftarrow Ah_{l}$
12:     $\alpha_{l+1}\leftarrow\frac{r_{l}^{T}z_{l}}{h_{l}^{T}v_{l}}$
13:     $u_{l+1}\leftarrow u_{l}+\alpha_{l+1}h_{l}$
14:     $r_{l+1}\leftarrow r_{l}-\alpha_{l+1}v_{l}$
15:     if $||r_{l+1}||_{2}<$ tolerance then
16:         early-stopping $\leftarrow$ true
17:     end if
18:     $z_{l+1}\leftarrow P^{-1}r_{l+1}$
19:     $\beta_{l+1}\leftarrow\frac{r_{l+1}^{T}z_{l+1}}{r_{l}^{T}z_{l}}$
20:     $h_{l+1}\leftarrow z_{l+1}+\beta_{l+1}h_{l}$
21:     $\tilde{T}_{l+1,l+1}\leftarrow\frac{1}{\alpha_{l+1}}+\frac{\beta_{l}}{\alpha_{l}}$
22:     if $l>0$ then
23:         $\tilde{T}_{l,l+1},\tilde{T}_{l+1,l}\leftarrow\frac{\sqrt{\beta_{l}}}{\alpha_{l}}$
24:     end if
25:     if early-stopping then
26:         return $u_{l+1},\tilde{T}$
27:     end if
28:end for
A.7 Zero fill-in incomplete Cholesky factorization
Algorithm 5 Zero fill-in incomplete Cholesky algorithm
1:Matrix $A\in\mathbb{R}^{m\times m}$, matrix $S\in\mathbb{R}^{m\times m}$ with sparsity pattern
2:Sparse lower triangular matrix $L$ with $A\approx LL^{T}$
3:for $i\leftarrow 1$ to $m$ do
4:     for $j\leftarrow 1$ to $m$ do
5:         if $(i,j)\in S$ and $i\geq j$ then
6:              $s\leftarrow L_{i\cdot}L_{j\cdot}^{T}$
7:              if $i==j$ then
8:                  $L_{ii}\leftarrow\sqrt{A_{ii}-s}$
9:              else
10:                  $L_{ij}\leftarrow\frac{A_{ij}-s}{L_{jj}}$
11:              end if
12:         end if
13:     end for
14:end for
A.8 Additional results for preconditioner comparison
Figure 9: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when the response variable follows a Bernoulli likelihood with a logit link function. The dashed line represents the result for the Cholesky decomposition.
Figure 10: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when $m_{2}=\frac{m_{1}}{2}$, $m=m_{1}+m_{2}=99^{\prime}999$, $n=10m$, and the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
Figure 11: Negative log-marginal likelihood and wall-clock time in seconds for different preconditioners and numbers of random vectors $t$ when $\sigma^{2}=0.25$, $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, and the response variable follows a Gaussian likelihood. The dashed line represents the result for the Cholesky decomposition.
A.9 Additional results for parameter estimation and prediction in simulated experiments
Krylov (GPBoost) Cholesky (GPBoost) lme4 glmmTMB
RMSE $\sigma^{2}$ $1.75\times 10^{-3}$ $1.75\times 10^{-3}$ $1.73\times 10^{-3}$ $1.73\times 10^{-3}$
Bias $\sigma^{2}$ $1.24\times 10^{-4}$ $1.26\times 10^{-4}$ $1.23\times 10^{-4}$ $1.23\times 10^{-4}$
RMSE $\sigma_{1}^{2}$ $8.97\times 10^{-3}$ $8.96\times 10^{-3}$ $8.83\times 10^{-3}$ $8.83\times 10^{-3}$
Bias $\sigma_{1}^{2}$ $-5.96\times 10^{-4}$ $-6.15\times 10^{-4}$ $-5.57\times 10^{-4}$ $-5.57\times 10^{-4}$
RMSE $\sigma_{2}^{2}$ $8.34\times 10^{-3}$ $8.35\times 10^{-3}$ $8.13\times 10^{-3}$ $8.13\times 10^{-3}$
Bias $\sigma_{2}^{2}$ $1.79\times 10^{-4}$ $1.56\times 10^{-4}$ $1.53\times 10^{-4}$ $1.54\times 10^{-4}$
RMSE $\beta_{0}$ $1.46\times 10^{-2}$ $1.46\times 10^{-2}$ $1.46\times 10^{-2}$ $1.46\times 10^{-2}$
Bias $\beta_{0}$ $1.36\times 10^{-3}$ $1.36\times 10^{-3}$ $1.36\times 10^{-3}$ $1.36\times 10^{-3}$
RMSE $\beta_{1}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$ $1.17\times 10^{-2}$
Bias $\beta_{1}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$ $-2.06\times 10^{-4}$
RMSE $\beta_{2}$ $1.08\times 10^{-2}$ $1.08\times 10^{-2}$ $1.08\times 10^{-2}$ $1.08\times 10^{-2}$
Bias $\beta_{2}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$ $-4.41\times 10^{-4}$
RMSE $\beta_{3}$ $1.25\times 10^{-2}$ $1.25\times 10^{-2}$ $1.25\times 10^{-2}$ $1.25\times 10^{-2}$
Bias $\beta_{3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$ $-1.14\times 10^{-3}$
RMSE $\beta_{4}$ $1.31\times 10^{-2}$ $1.31\times 10^{-2}$ $1.31\times 10^{-2}$ $1.31\times 10^{-2}$
Bias $\beta_{4}$ $1.28\times 10^{-3}$ $1.28\times 10^{-3}$ $1.28\times 10^{-3}$ $1.28\times 10^{-3}$
RMSE $\beta_{5}$ $1.06\times 10^{-2}$ $1.06\times 10^{-2}$ $1.06\times 10^{-2}$ $1.06\times 10^{-2}$
Bias $\beta_{5}$ $3.21\times 10^{-4}$ $3.21\times 10^{-4}$ $3.21\times 10^{-4}$ $3.21\times 10^{-4}$
Table 2: Root mean squared error (RMSE) and bias of the variance and coefficient estimates for Gaussian likelihoods.
Krylov (GPBoost) Cholesky (GPBoost) glmmTMB
RMSE $\sigma_{1}^{2}$ $2.01\times 10^{-2}$ $2.01\times 10^{-2}$ $2.05\times 10^{-2}$
Bias $\sigma_{1}^{2}$ $-1.15\times 10^{-2}$ $-1.15\times 10^{-2}$ $-1.13\times 10^{-2}$
RMSE $\sigma_{2}^{2}$ $2.15\times 10^{-2}$ $2.15\times 10^{-2}$ $2.20\times 10^{-2}$
Bias $\sigma_{2}^{2}$ $-1.36\times 10^{-2}$ $-1.36\times 10^{-2}$ $-1.36\times 10^{-2}$
RMSE $\beta_{0}$ $1.73\times 10^{-2}$ $1.73\times 10^{-2}$ $1.73\times 10^{-2}$
Bias $\beta_{0}$ $-3.68\times 10^{-4}$ $-3.67\times 10^{-4}$ $-4.76\times 10^{-4}$
RMSE $\beta_{1}$ $5.18\times 10^{-2}$ $5.18\times 10^{-2}$ $5.10\times 10^{-2}$
Bias $\beta_{1}$ $1.33\times 10^{-3}$ $1.34\times 10^{-3}$ $7.09\times 10^{-4}$
RMSE $\beta_{2}$ $5.33\times 10^{-2}$ $5.33\times 10^{-2}$ $5.19\times 10^{-2}$
Bias $\beta_{2}$ $-6.51\times 10^{-3}$ $-6.53\times 10^{-3}$ $-5.87\times 10^{-3}$
RMSE $\beta_{3}$ $5.10\times 10^{-2}$ $5.09\times 10^{-2}$ $5.05\times 10^{-2}$
Bias $\beta_{3}$ $6.00\times 10^{-4}$ $5.82\times 10^{-4}$ $1.86\times 10^{-4}$
RMSE $\beta_{4}$ $5.25\times 10^{-2}$ $5.25\times 10^{-2}$ $5.24\times 10^{-2}$
Bias $\beta_{4}$ $-1.24\times 10^{-3}$ $-1.23\times 10^{-3}$ $-1.30\times 10^{-3}$
RMSE $\beta_{5}$ $5.18\times 10^{-2}$ $5.18\times 10^{-2}$ $4.99\times 10^{-2}$
Bias $\beta_{5}$ $-6.36\times 10^{-3}$ $-6.36\times 10^{-3}$ $-5.61\times 10^{-3}$
Table 3: Root mean squared error (RMSE) and bias of the variance and coefficient estimates for Bernoulli likelihoods.
Krylov (GPBoost) Cholesky (GPBoost) lme4 glmmTMB
$\overline{\text{RMSE}}$ $1.62\times 10^{-1}$ $1.62\times 10^{-1}$ $1.62\times 10^{-1}$ $1.62\times 10^{-1}$
sd($\overline{\text{RMSE}}$) $2.32\times 10^{-4}$ $2.32\times 10^{-4}$ $2.32\times 10^{-4}$ $2.32\times 10^{-4}$
$\overline{\text{LS}}$ $-4.01\times 10^{-1}$ $-4.01\times 10^{-1}$
sd($\overline{\text{LS}}$) $1.40\times 10^{-3}$ $1.40\times 10^{-3}$
Table 4: Average root mean squared error (RMSE) for predictive means and average log score (LS) for probabilistic predictions with corresponding standard errors for Gaussian likelihoods.
Krylov (GPBoost) Cholesky (GPBoost) glmmTMB
$\overline{\text{RMSE}}$ $5.00\times 10^{-1}$ $5.00\times 10^{-1}$ $5.00\times 10^{-1}$
sd($\overline{\text{RMSE}}$) $5.37\times 10^{-4}$ $5.37\times 10^{-4}$ $5.37\times 10^{-4}$
$\overline{\text{LS}}$ $7.26\times 10^{-1}$ $7.26\times 10^{-1}$
sd($\overline{\text{LS}}$) $1.12\times 10^{-3}$ $1.12\times 10^{-3}$
Table 5: Average root mean squared error (RMSE) for predictive means and average log score (LS) for probabilistic predictions with corresponding standard errors for Bernoulli likelihoods.
Figure 12: Average wall clock times (s) for parameter estimation and different $m$ with $m_{2}=\frac{m_{1}}{2}$. The dimensions of the random effects $m$ are chosen so that they are divisible by three and correspond as closely as possible to those in Figure 6. Simulated data follows either a Gaussian or a Bernoulli likelihood.
A.10 Additional results for the real-world applications
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 13: Non-zero entries of the matrix $Z^{T}Z$ for the real-world data sets.
data set method Time (s) nll_optimum
cars Cholesky (GPBoost) 104.3 27518.0
Krylov (GPBoost) 7.7 27517.1
glmmTMB 1152.2 27518.1
lme4 62.3 27518.0
building_permits Cholesky (GPBoost) 5647.5 867567.7
Krylov (GPBoost) 97.4 867566.5
glmmTMB 3763.2 867567.8
lme4 600.0 867567.3
instEval Cholesky (GPBoost) 9.4 118764.0
Krylov (GPBoost) 0.5 118763.5
glmmTMB 76.1 118764.0
lme4 11.5 118764.0
employee_access Cholesky (GPBoost) 628.0 5494.4
Krylov (GPBoost) 80.5 5492.3
glmmTMB 287.9 5494.4
lme4 8362.9 5504.5
upselling Cholesky (GPBoost) 1747.8 11488.5
Krylov (GPBoost) 41.1 11488.4
Table 6: Time for parameter estimation and negative log-marginal likelihood at the optimum for different real-world data sets and models.
data set method $\sigma^{2}$ $\sigma_{1}^{2}$ $\sigma_{2}^{2}$ $\sigma_{3}^{2}$
cars Cholesky (GPBoost) $5.24\times 10^{-2}$ $3.28\times 10^{-1}$ $2.04\times 10^{-1}$
Krylov (GPBoost) $5.24\times 10^{-2}$ $3.28\times 10^{-1}$ $2.04\times 10^{-1}$
glmmTMB $5.24\times 10^{-2}$ $3.28\times 10^{-1}$ $2.04\times 10^{-1}$
lme4 $5.24\times 10^{-2}$ $3.28\times 10^{-1}$ $2.04\times 10^{-1}$
building_permits Cholesky (GPBoost) $1.31\times 10^{0}$ $9.10\times 10^{-1}$ $3.98\times 10^{-2}$ $3.97\times 10^{-2}$
Krylov (GPBoost) $1.31\times 10^{0}$ $9.09\times 10^{-1}$ $3.98\times 10^{-2}$ $3.96\times 10^{-2}$
glmmTMB $1.31\times 10^{0}$ $9.10\times 10^{-1}$ $3.69\times 10^{-2}$ $4.25\times 10^{-2}$
lme4 $1.31\times 10^{0}$ $9.10\times 10^{-1}$ $7.93\times 10^{-2}$ $1.08\times 10^{-4}$
instEval Cholesky (GPBoost) $1.38\times 10^{0}$ $1.07\times 10^{-1}$ $2.57\times 10^{-1}$
Krylov (GPBoost) $1.38\times 10^{0}$ $1.07\times 10^{-1}$ $2.57\times 10^{-1}$
glmmTMB $1.38\times 10^{0}$ $1.07\times 10^{-1}$ $2.57\times 10^{-1}$
lme4 $1.38\times 10^{0}$ $1.07\times 10^{-1}$ $2.57\times 10^{-1}$
employee_access Cholesky (GPBoost) $1.52\times 10^{0}$ $5.78\times 10^{0}$ $9.48\times 10^{-6}$
Krylov (GPBoost) $1.52\times 10^{0}$ $5.79\times 10^{0}$ $1.17\times 10^{-5}$
glmmTMB $1.52\times 10^{0}$ $5.78\times 10^{0}$ $3.14\times 10^{-10}$
lme4 $1.60\times 10^{0}$ $6.06\times 10^{0}$ $0.00\times 10^{0}$
upselling Cholesky (GPBoost) $3.56\times 10^{-1}$ $2.59\times 10^{-1}$ $9.86\times 10^{-3}$
Krylov (GPBoost) $3.58\times 10^{-1}$ $2.58\times 10^{-1}$ $9.58\times 10^{-3}$
Table 7: Estimates for $\sigma^{2}$, $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, and $\sigma_{2}^{3}$ for different real-world data sets and models. For reasons of space, we do not report estimates for other variance parameters.
data set method $\beta_{0}$ $\beta_{1}$ $\beta_{2}$ $\beta_{3}$
cars Cholesky (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
Krylov (GPBoost) $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
glmmTMB $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.63\times 10^{-3}$
lme4 $1.02\times 10^{1}$ $3.50\times 10^{-1}$ $-1.77\times 10^{-1}$ $-2.64\times 10^{-3}$
building_permits Cholesky (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.01\times 10^{-3}$
Krylov (GPBoost) $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.00\times 10^{-3}$
glmmTMB $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.02\times 10^{-3}$
lme4 $3.62\times 10^{0}$ $5.44\times 10^{-4}$ $-1.36\times 10^{-5}$ $2.01\times 10^{-3}$
instEval Cholesky (GPBoost) $3.31\times 10^{0}$ $5.21\times 10^{-2}$ $7.23\times 10^{-2}$ $1.37\times 10^{-1}$
Krylov (GPBoost) $3.31\times 10^{0}$ $5.21\times 10^{-2}$ $7.23\times 10^{-2}$ $1.37\times 10^{-1}$
glmmTMB $3.31\times 10^{0}$ $5.21\times 10^{-2}$ $7.23\times 10^{-2}$ $1.37\times 10^{-1}$
lme4 $3.31\times 10^{0}$ $5.21\times 10^{-2}$ $7.23\times 10^{-2}$ $1.37\times 10^{-1}$
employee_access Cholesky (GPBoost) $7.21\times 10^{0}$
Krylov (GPBoost) $7.16\times 10^{0}$
glmmTMB $7.19\times 10^{0}$
lme4 $6.92\times 10^{0}$
upselling Cholesky (GPBoost) $-1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.42\times 10^{-3}$ $3.86\times 10^{-6}$
Krylov (GPBoost) $-1.74\times 10^{0}$ $2.32\times 10^{-5}$ $9.48\times 10^{-3}$ $3.94\times 10^{-6}$
Table 8: Estimates for $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ for different real-world data sets and models. For reasons of space, we do not report estimates for other coefficients.